### v3

# Rainbow triangles and the Erdős-Hajnal problem in projective geometries

Abstract.
We formulate a geometric version of the Erdős-Hajnal conjecture that applies to finite projective geometries rather than graphs, in both its usual ‘induced’ form and the multicoloured form. The multicoloured conjecture states, roughly, that a colouring $c$ of the points of $\operatorname{PG}(n-1,q)$ containing no copy of a fixed colouring $c_{0}$ of $\operatorname{PG}(k-1,q)$ for small $k$ must contain a subspace of dimension polynomial in $n$ that avoids some colour.
If $(k,q)=(2,2)$, then $c_{0}$ is a colouring of a three-element ‘triangle’, and there are three essentially different cases, all of which we resolve. We derive both the cases where $c_{0}$ assigns the same colour to two different elements from a recent breakthrough result in additive combinatorics due to Kelley and Meka. We handle the case that $c_{0}$ is a ‘rainbow’ colouring by proving that rainbow-triangle-free colourings of projective geometries are exactly those that admit a certain decomposition into two-coloured pieces. This is closely analogous to a theorem of Gallai on rainbow-triangle-free coloured complete graphs.
We also show that existing structure theorems resolve certain two-coloured cases where $(k,q)=(2,3)$, and $(k,q)=(3,2)$.
Key words and phrases:
matroids, graphs, colourings
1991 Mathematics Subject Classification:
05B35
1. Introduction
The Erdős-Hajnal conjecture [9] states that for every graph $H$, there is a constant $\beta=\beta(H)>0$ such that, if $G$ is a graph on $n$ vertices with no induced $H$-subgraph, then $G$ has a clique or stable set of size at least $n^{\beta}$. This problem is known for some particular graphs $H$ (notably the $5$-cycle) as well as when $H$ belongs to one of a few infinite classes such as the cographs, but is wide open in general. See [8] for a survey.
Erdős and Hajnal [10] in fact stated a multicoloured generalization of the same conjecture. To state this, we introduce some terminology. A colouring of a finite set $X$ is any function $c$ whose domain is $X$, and, for $s\in\mathbb{N}$, an $s$-colouring is a function $c:X\to\{1,\dotsc,s\}$. Given a colouring $c$, we write $[c]$ for the range of $c$ (the set of colours used by $c$), and $|c|$ for the number of colours used by $c$.
For a graph $G$, a colouring of $G$ means a colouring of $E(G)$. The conjecture is stated in terms of colourings of complete graphs ‘containing’ other colourings. If $c$ and $d$ are colourings of complete graphs $H$ and $G$ respectively, then we say that $d$ contains $c$ if there is an injection $\psi:V(H)\to V(G)$ such that $c(xy)=d(\psi(x)\psi(y))$ for all distinct $x,y\in V(H)$. A colouring not containing $c$ is $c$-free.
Conjecture 1.1 (Multicoloured Erdős-Hajnal conjecture for graphs).
For all $1\leq s_{0}\leq s$ and $k\geq 1$, and every $s_{0}$-colouring $c_{0}$ of $K_{k}$, there exist $\delta,C>0$ such that, for all $n\in\mathbb{N}$ and every $c_{0}$-free $s$-colouring $c$ of $K_{n}$, there is a colour $i\in\{1,\dotsc,s\}$ and a set $X$ of at least $Cn^{\delta}$ vertices of $K_{n}$ such that $c(e)\neq i$ for every edge $e$ with both ends in $X$.
The set $X$ is referred to as homogeneous. The case where $s_{0}=s=2$ is the classical Erdős-Hajnal conjecture.
We make an analogous conjecture about colourings of projective geometries. In this context, we write $G\cong\operatorname{PG}(n-1,q)$ to denote the projective space associated with some generic $n$-dimensional vector space $V$ over $\operatorname{GF}(q)$; the points of $G$ are the one-dimensional subspaces of $V$. Going forward, we often identify this object with the corresponding simple rank-$n$ $\operatorname{GF}(q)$-representable matroid on $\tfrac{q^{n}-1}{q-1}$ elements.
A colouring of $G$ is a colouring of the $\tfrac{q^{n}-1}{q-1}$ points of $G$. If $c$ and $d$ are colourings of $H\cong\operatorname{PG}(m-1,q)$ and $G\cong\operatorname{PG}(n-1,q)$ respectively, then we say that $d$ contains $c$ if there is an injective homomorphism $\psi:H\to G$ such that $c(x)=d(\psi(x))$ for every point $x$ of $H$. If $d$ does not contain $c$, then $d$ is $c$-free. We are primarily interested in the binary case, where one can think of $n$-dimensional binary projective space more concretely as the set of nonzero vectors in $\mathbb{F}_{2}^{n}$, and homomorphisms as linear maps.
Given $s\in\mathbb{N}$, and an $s$-colouring $c$ of $G\cong\operatorname{PG}(n-1,2)$, we say that a subspace $V$ of $G$ is $(c,s)$-homogeneous if there is some $i\in\{1,\dotsc,s\}$ for which $i\notin c(V)$. That is, the colouring $c$ avoids some colour on $V$.
Conjecture 1.2 (Multicoloured geometric Erdős-Hajnal conjecture).
For every prime power $q$, all $k,s_{0},s\in\mathbb{N}$ with $s_{0}\leq s$, and every $s_{0}$-colouring $c_{0}$ of $\operatorname{PG}(k-1,q)$, there exists $\delta,C>0$ such that, for all $n\in\mathbb{N}$ and every $c_{0}$-free $s$-colouring $c$ of $G\cong\operatorname{PG}(n-1,q)$, there is a $(c,s)$-homogeneous subspace of $G$ with dimension at least $Cn^{\delta}$.
Induced restrictions
The classical Erdős-Hajnal conjecture on induced subgraphs is precisely the case of Conjecture 1.1 where both the colourings use only two colours. Specializing Conjecture 1.2 in the same way gives the following conjecture, phrased in terms of induced restrictions. In what follows, we slightly abuse notation by identifying a projective geometry $\operatorname{PG}(n-1,q)$ with its set of points. If $H$ and $X$ are sets of points in $\operatorname{PG}(k-1,q)$ and $\operatorname{PG}(n-1,q)$ respectively, then we say that $X$ contains $H$ as an induced restriction if there is an injective homomorphism $\psi:\operatorname{PG}(k-1,q)\to\operatorname{PG}(n-1,q)$ such that, for each point $x$ of $\operatorname{PG}(k-1,q)$, we have $\psi(x)\in X$ if and only if $x\in H$.
Conjecture 1.3 (Geometric Erdős-Hajnal conjecture).
For every prime power $q$ and all $k\in\mathbb{N}$ and $H\subseteq\operatorname{PG}(k-1,q)$, there exists $\delta,C>0$ such that, for all $n\in\mathbb{N}$ and $X\subseteq\operatorname{PG}(n-1,q)$ not containing $H$ as an induced restriction, there is a subspace $W$ of $\operatorname{PG}(n-1,q)$ such that $\dim(W)>Cn^{\delta}$, and either $W\subseteq X$ or $W\cap X=\varnothing$.
Finally, we state a strengthening of the last two conjectures, suggested by Jacob Fox, which asserts that in fact they hold for linear functions rather than polynomial. This may seem quite strong, but we know of no construction that refutes it.
Conjecture 1.4.
Conjectures 1.2 and 1.3 hold with $\delta=1$.
Results
[…]
To prove Conjecture 1.2, it suffices to consider the case where the colouring $c_{0}$ is surjective: that is, where $s_{0}=|c_{0}|$. Thus, by symmetry, there are only three essentially different colourings $c_{0}$ of $T=\{x_{1},x_{2},x_{3}\}$: the $1$-colouring $c_{0}$ with $c_{0}]=\{1\}$, the $2$-colouring $c_{0}$ where $c_{0}(x_{1})=c_{0}(x_{2})=1$ and $c_{0}(x_{3})=2$, and the ‘rainbow’ $3$-colouring with $c_{0}(x_{i})=i$ for all $i$. Our first result handles the first two cases, for arbitrary $s\geq s_{0}$.
Theorem 1.5.
Let $s_{0}\in[2]$ and $s\geq s_{0}$, and $c_{0}$ be an $s_{0}$-colouring of $T$. Then Conjecture 1.2 holds for $c_{0}$ and $s$, with $\delta=\tfrac{1}{7}$.
The proof of this theorem is via a reduction to a recent breakthrough result in additive combinatorics due to Kelley and Meka [16], formulated by Hunter and Pohoata [15] in our setting.
Our next result handles the rainbow colouring of $T$, when $s_{0}=s=3$.
Theorem 1.6.
Let $c_{0}$ be the rainbow $3$-colouring of $T$. Then Conjecture 1.2 holds for $c_{0}$ and $s=3$, with $C=\tfrac{1}{3}$ and $\delta=1$. Moreover, the values of $\delta$ and $C$ are best-possible.
The proof is via very different techniques. Following a route analogous to results due to Gallai [12, 18] and Fox, Grinshpun and Pach [11], we prove an exact structure theorem, Theorem 6.7, which describes the $3$-colourings of $\operatorname{PG}(n-1,2)$ with no rainbow triangle, and derive Theorem 1.6 as an easy consequence.
Theorem 6.7 is a little too technical to state precisely here, but it roughly states that, if $c$ is a colouring of the points of a projective geometry having no three colinear points of different colours, then $c$ can be constructed from two-coloured subspaces using a certain ‘join’ operation. The following weakening will be enough to imply Theorem 1.6.
Theorem 1.7.
Let $n\in\mathbb{N}$, let $q$ be a prime power, and $c$ be a colouring of the points of $G\cong\operatorname{PG}(n-1,q)$ so that $|c(L)|\leq 2$ for each two-dimensional subspace $L$ of $G$. Then there are subspaces $\varnothing=W_{0}\subseteq W_{1}\subseteq\dotsc\subseteq W_{k}=G$ of $G$ such that $|c(W_{i+1}-W_{i})|\leq 2$ for all $0\leq i<k$.
Note that this applies to all finite projective geometries, not just those over $\mathbb{F}_{2}$; this is also the case for Theorem 6.7. Sadly, for larger fields, Theorem 6.7 has no direct applications to Conjecture 1.2, because the notion of a ‘line’ diverges from that of a ‘triangle’.
In the case where the number of colours is equal to the dimension, our structure theorem has an especially nice consequence, stating that there is a unique $n$-colouring of $\operatorname{PG}(n-1,q)$ with no rainbow triangle, up to automorphisms and permutations of the colour set. In the case where $q=2$, this follows from a much more general result of Bérczi and Schwarcz [2, Theorem 4], who call these colourings ‘standard’. (Their result has an apparently stronger assumption that omits rainbow circuits rather than triangles, but we will see in Theorem 5.2 that this hypothesis is equivalent to ours in the case of projective geometries.)
Theorem 1.8.
Let $q$ be a prime power, and $c$ be an $n$-colouring of $\operatorname{PG}(n-1,q)$ that uses all $n$ colours, such that each line contains at most two colours. Then there is a maximal chain of subspaces $\varnothing=F_{0}\subset\dotsc\subset F_{n}=\operatorname{PG}(n-1,q)$ such that the sets $F_{i}-F_{i-1}$ are monochromatic with different colours.
Conjecture 1.3 is equivalent to the case of Conjecture 1.2 where $s=s_{0}=2$, so is resolved in the case where $q=k=2$ by Theorem 1.5. We can resolve two more cases of Conjecture 1.5 (up to complementation) using known exact structure theorems.
The first is in the case where $q=3$ and $k=2$, and $H$ is a set of two distinct elements in $\operatorname{PG}(1,3)$. Viewed as a two-colouring, this object $H$ is self-complementary, since $\operatorname{PG}(1,3)$ has four elements and symmetric automorphism group. Conjecture 1.3 will follow for $H$ from a characterization of the $H$-free matroids due to Mizell and Oxley [20], which we prove separately and in more generality as Theorem 7.9.
Theorem 1.9.
Let $H\subseteq\operatorname{PG}(1,3)$ with $|H|=2$. Then Conjecture 1.3 holds for $H$, with $\delta=1$ and $C=\tfrac{1}{2}$. These values are best-possible.
The second case we can settle is that of the claw, which is defined as a basis of the seven-point Fano plane $\operatorname{PG}(2,2)$. This will follow from the main theorem of [21], which allows us to reduce the problem to Theorem 1.5.
Theorem 1.10.
Let $H\subseteq\operatorname{PG}(2,2)$ consist of either a (three-element) basis of $\operatorname{PG}(2,2)$ or its four-element complement. Then Conjecture 1.3 holds for $H$, with $\delta=\tfrac{1}{7}$.
2. Preliminaries
We write $\mathbb{N}$ for the set $\{0,1,\dotsc\}$, and $n]=\{1,\dotsc,n\}$ for each $n\in\mathbb{N}$. When discussing matroids, we usually use the notation and terminology of Oxley [22]. For sets $X,Y$ in a matroid $M$, we write $\sqcap_{M}(X,Y)$ for the ‘local connectivity’ evaluated as $r_{M}(X)+r_{M}(Y)-r_{M}(X\cup Y)$. A rank-one flat is a point, and a rank-two flat is a line. Two flats $F,F^{\prime}$ of a matroid $M$ are complementary if $\sqcap_{M}(F,F^{\prime})=0$ and $F\cup F^{\prime}$ is spanning in $M$; such an $F^{\prime}$ is a complement of $F$, and satisfies $r(M)=r_{M}(F)+r_{M}(F^{\prime})$. Note that if $B$ is any basis for $M$ containing a basis $I$ for the flat $F$, then $\operatorname{cl}_{M}(B-I)$ is an example of a complement of $F$, so every flat has at least one complement.
Our first tool is due to Bose and Burton [3]. A projective subspace of $\mathbb{F}_{2}^{n}$ is a linear subspace with the zero vector removed.
Theorem 2.1.
[…]
Corollary 2.2.
Each $X\subseteq\mathbb{F}_{2}^{n}$ satisfies $|X|\geq 2^{n-\alpha(X)}-1$.
Finally, the qualitative Ramsey result that facilitates the discussion ahead is due to Graham, Leeb and Rothschild [13].
Theorem 2.3.
There is a function $R_{q}(s,t)$ so that, for each prime power $q$ and all $s,t\in\mathbb{N}$, for all $n\in\mathbb{N}$ with $n\geq R_{q}(s,t)$ and each colouring $c$ of $G\cong\operatorname{PG}(n-1,q)$ with $c]\subseteq\{\mathsf{Red},\mathsf{Blue}\}$, there is either an $s$-dimensional subspace of $G$ coloured $\mathsf{Red}$, or a $t$-dimensional subspace of $G$ coloured $\mathsf{Blue}$.
3. Monochromatic Triangles
In this section, we prove Theorem 1.5. We first review the relevant material in additive combinatorics, and state the main theorem of [15] in a slightly stronger form. Logarithms here are all binary, and for $\subseteq\mathbb{F}_{2}^{n}, X+YX+Y$ denotes the sum-set $\{x+y\colon x\in X,y\in Y\}$.
For $\delta\in(0,1]$, we follow [15] by writing $\mathcal{L}(\delta)=\lceil\log(1/\delta)\rceil$ for the smallest integer $k$ such that $\delta 2^{k}\geq 1$; note that $k\leq n$. For $X,V\subseteq\mathbb{F}_{2}^{n}$, write $\mu_{V}(X)=\frac{|X\cap V|}{|V|}$ and $\mu(X)=\mu_{\mathbb{F}_{2}^{n}}(X)=\frac{|X|}{2^{n}}$. The following is a convenient but slightly weaker rephrasing of Theorem 2.1 in this language.
Theorem 3.1.
If $\varnothing\neq X\subseteq\mathbb{F}_{2}^{n}$, then $\alpha(X)+1\geq\mathcal{L}(\mu(X))$.
Proof.
We may assume that $\alpha(X)<n$. By Corollary 2.2, we have
$\mathcal{L}(\mu(X))=\left\lceil\log\left(\tfrac{2^{n}}{|X|}\right)\right\rceil=n-\left\lfloor\log|X|\right\rfloor\leq n-\left\lfloor\log(2^{n-\alpha(X)}-1)\right\rfloor.$
Since $\alpha(X)<n$, we have $\left\lfloor\log(2^{n-\alpha(X)}-1)\right\rfloor=n-\alpha(X)-1$, and the result follows. ∎
The following is the hard result we need from additive combinatorics. It follows from a specialized strengthening to $\mathbb{F}_{2}$, due to Hunter and Pohoata, of recent work of Kelley and Meka [16]. The statement below is a weakening of [15, Proposition 3.1] in the case where $r=1$.
Theorem 3.2.
Let $A\subseteq\mathbb{F}_{2}^{n}$, and $\delta,\gamma\in(0,1)$. There exists a subspace $V$ of $\mathbb{F}_{2}^{n}$ with $\operatorname{codim}(V)=O(\mathcal{L}(\delta)^{5}\mathcal{L}(\gamma)^{2})$ such that either $\mu_{V}(A)<\delta$, or $\mu_{V}(A+A)>1-\gamma$.
Using this, we can prove a key result showing that for every set $X\subseteq\mathbb{F}_{2}^{n}$, either the sum-set $X+X$ or the complement of $X$ contains a subspace whose dimension is polynomial in $n$. The proof is a straightforward consequence of the techniques in [15].
Theorem 3.3.
[…]
Proof.
Let $C^{\prime}$ be the constant implicit in Lemma 3.2. We show that $C=\min(1-2^{-7},\tfrac{1}{4}(C^{\prime})^{-1/7})$ satisfies the theorem. Let $X\subseteq\mathbb{F}_{2}^{n}$ and let $d=\lceil Cn^{1/7}\rceil$. If $Cn^{1/7}\leq 1$, the result is trivial, so we may assume that $d\leq 2Cn^{1/7}$.
Let $V$ be a subspace given by Lemma 3.2 with $\delta=\gamma=2^{-d}$, so $\mathcal{L}(\delta)=\mathcal{L}(\gamma)=d$ and thus, by the choice of $C$, we have
$\dim(V)\geq n-C^{\prime}d^{7}\geq n-C^{\prime}(2Cn^{1/7})^{7}\geq n(1-2^{-7})\geq Cn^{1/7}.$
By the choice of $\delta,\gamma$ and $V$, there is a set $Y$ equal to one of $X\cap V$ or $V\backslash(X+X)$ for which $\mu_{V}(Y)<2^{-d}$. If $Y=X\cap V$ then $\alpha(X)\geq\alpha_{V}(Y)$ and if $Y=V\backslash(X+X)$, then $\omega(X+X)\geq\omega((X+X)\cap V)=\alpha(V\backslash(X+X)))=\alpha_{V}(Y)$, so $\max(\alpha(X),\omega(X+X))\geq\alpha_{V}(Y)$ in both cases, and it suffices to show that $\alpha_{V}(Y)\geq Cn^{1/7}$.
If $Y=\varnothing$, then $\alpha_{V}(Y)=\dim(V)\geq Cn^{1/7}$ as required. Otherwise, since $\mu_{Y}(V)<2^{-d}$ we have $\mathcal{L}(\mu_{V}(Y))\geq d+1$, so Theorem 3.1 gives $\alpha_{V}(Y)+1\geq\mathcal{L}(\mu_{V}(Y))\geq d+1\geq Cn^{1/7}+1$ and so $\alpha_{V}(Y)\geq Cn^{1/7}$. ∎
This easily implies the following estimate on the Ramsey number $R_{2}(2,t)$ as defined in Theorem 2.3; this was also stated in [15].
Theorem 3.4.
There is a constant $C$ so that $R_{2}(2,t)\leq Ct^{7}$ for all $t$.
Proof of Theorem 3.4.
[…]
Analogously to [15, Conjecture 1], we conjecture that the right bound in Theorem 3.3 is linear.
Conjecture 3.5.
There is a constant $C>0$ such that, for all $n\in\mathbb{N}$, every $X\subseteq\mathbb{F}_{2}^{n}$ satisfies $\max(\alpha(X),\omega(X+X))\geq Cn$.
We can now prove Theorem 1.5.
Theorem 3.6.
[…]
Proof.
[…]
If $\alpha(X)\geq Cn^{1/7}$, then there is a subspace $W$ of $\mathbb{F}_{2}^{n}$ containing no point coloured $\kappa$, for which $\dim(W)\geq Cn^{1/7}$. Thus, $W$ gives a $(c,s)$-homogeneous subspace of the same dimension in $G$.
[…]
4. Modularity
Our main results will be applied to projective geometries over finite fields, but it is convenient to work at the more abstract level of modular matroids. The consequence is that our proofs are co-ordinate free, and apply to a slightly larger class of objects.
A matroid $M$ is modular if every pair $F,F^{\prime}$ of flats of $M$ satisfies $r_{M}(F)+r_{M}(F^{\prime})=r_{M}(F\cup F^{\prime})+r_{M}(F\cap F^{\prime})$, or equivalently if $r_{M}(\operatorname{cl}_{M}(A)\cap\operatorname{cl}_{M}(B))=\sqcap_{M}(A,B)$ for all $A,B\subseteq E(M)$.
Canonical examples include matroids of rank at most two, free matroids, projective planes in rank three (including non-Desarguesian ones), and projective geometries over finite fields. Indeed, the simple connected modular matroids of rank at least four are precisely those of the form $\operatorname{PG}(n-1,q)$ for a prime power $q$ (See [22, Proposition 6.9.1]).
Even though Theorem 1.6 only applies to binary projective geometries, Theorem 6.7 applies to all connected modular matroids, which by the aforementioned classification comprise precisely the loopless matroids whose simplification is either a rank-$2$ uniform matroid, a projective plane, or $\operatorname{PG}(n-1,q)$ for $n\geq 4$.
The first four lemmas in this section follow easily from the classification just discussed, but for completeness, we include self-contained proofs that use the definition of modularity directly.
Lemma 4.1.
If $M$ is a loopless modular matroid, then flats $F,F^{\prime}$ of $M$ are complementary if and only if they are disjoint and their union spans $M$.
Proof.
It suffices to show that two flats $F_{1},F_{2}$ have local connectivity zero if and only if they are disjoint. But $r_{M}(F_{1}\cap F_{2})=\sqcap_{M}(F_{1},F_{2})$, and since $M$ is loopless, we have $r_{M}(F_{1}\cap F_{2})=0$ if and only if $|F_{1}\cap F_{2}|=0$. The equivalence holds. ∎
Lemma 4.2.
If $M$ is a modular matroid, then $M/C$ is modular for each set $C\subseteq E(M)$, and $M|F$ is modular for each flat $F$ of $M$.
Proof.
The fact that $M|F$ is modular follows immediately from the definition, since the flats of $M|F$ are the flats of $M$ that are contained in $F$. To see that $M/C$ is modular for all $C$, let $X,Y\subseteq E(M/C)$. By the modularity of $M$ and the fact that $C\subseteq\operatorname{cl}_{M}(X\cup C)\cap\operatorname{cl}_{M}(Y\cup C)$,
$\sqcap_{M/C}(X,Y)$ $=\sqcap_{M}(X\cup C,Y\cup C)-r_{M}(C)$
$=r_{M}(\operatorname{cl}_{M}(X\cup C)\cap\operatorname{cl}_{M}(Y\cup C))-r_{M}(C)$
$=r_{M/C}(\operatorname{cl}_{M/C}(X)\cap\operatorname{cl}_{M/C}(Y)),$
so $M/C$ is modular. ∎
Lemma 4.3.
If $F$ is a nonempty proper flat of a connected modular matroid $M$, then $M|F$ and $M/F$ are both connected and modular.
Proof.
We know both are modular from the previous lemma, so it suffices to show that both are connected. To show that $M|F$ is connected, let $x,y\in F$. By the connectedness of $M$, there is a circuit $C$ of $M$ containing $x$ and $y$. If $C=\{x,y\}$, then $C$ is a circuit of $M|F$ containing $x$ and $y$, as required. Otherwise, both $\{x,y\}$ and $C-\{x,y\}$ are independent in $M$, so
$r_{M}(\operatorname{cl}_{M}(\{x,y\})\cap\operatorname{cl}_{M}(C-\{x,y\}))$ $=\sqcap_{M}(C-\{x,y\},\{x,y\})$
$=(|C|-2)+2-(|C|-1)=1.$
Therefore $M$ has a nonloop $e$ spanned by both $\{x,y\}$ and $C-\{x,y\}$. Since $F$ is a flat, $e\in F$. If $e$ is parallel to some $f\in\{x,y\}$, then the independent set $(C-\{x,y\})\cup\{f\}$ is spanned by its proper subset $C-\{x,y\}$, a contradiction. Therefore $e$ is parallel to neither $x$ nor $y$ but is spanned by $\{x,y\}$, so $\{e,x,y\}$ is a circuit contained in $F$. Therefore $M|F$ is connected.
To see that $M/F$ is connected, let $F_{0}$ be a complement of $F$, so $F_{0}$ is spanning in $M/F$. Since $F$ is a flat, the matroid $M/F$ is loopless and has a connected spanning restriction $M|F_{0}$, so is connected. ∎
The next lemma is a well-known result about finite projective planes, stated in matroidal language.
Lemma 4.4.
If $M$ is a connected, modular rank-three matroid, then there is an integer $q\geq 2$ such that every line of $M$ has length $q+1$, every point is contained in exactly $q+1$ lines, and the number of points and the number of lines are both equal to $q^{2}+q+1$.
Proof.
We may assume that $M$ is simple. By modularity, any two distinct lines intersect in a single element. Let $\mathcal{L}$ be the set of lines of $M$, and $L,L^{\prime}\in\mathcal{L}$ be distinct lines intersecting in a point $e$. Let $f\in L-e$ and $f^{\prime}\in L^{\prime}-e$, noting that $f\neq f^{\prime}$.
If $E(M)=L\cup L^{\prime}$, then since $M$ is coloopless, there are distinct $x,x^{\prime}\in L-L^{\prime}$ and distinct $y,y^{\prime}\in L^{\prime}-L$. Then $\operatorname{cl}_{M}(\{x,y\})$ and $\operatorname{cl}_{M}(\{x^{\prime},y^{\prime}\})$ intersect in some element $z$, and it is easy to argue that $z\notin L\cup L^{\prime}$, a contradiction. Therefore there exists $g\in E(M)-(L\cup L^{\prime})$.
Define $\varphi:L\to L^{\prime}$ by, for each $x\in L$, setting $\varphi(x)$ to be the unique point in $L^{\prime}\cap\operatorname{cl}_{M}(\{x,g\})$. Since $x$ is the only element of $L\cap\operatorname{cl}_{M}(\{\varphi(x),g^{\prime}\})$, the function $\varphi$ is injective. It follows that $|L^{\prime}|\geq|L|$. The choices of $L$ and $L^{\prime}$ were arbitrary, so all lines of $M$ have the same length $q+1$ for some $q\geq 1$.
Let $e\in E(M)$. Since $r(M)=3$, there is a line $L$ not containing $e$. The collection $\mathcal{L}_{e}=\{\operatorname{cl}_{M}(\{e,x\}):x\in L\}$ contains $|L|=q+1$ distinct lines, and since every line through $e$ intersects $L$, each line through $e$ belongs to $\mathcal{L}_{e}$. Therefore $e$ is in exactly $|\mathcal{L}_{e}|=q+1$ lines.
Moreover, the lines in $\mathcal{L}_{e}$ partition $E(M)-\{e\}$ and each has length $q+1$, so $|M|=1+\sum_{L\in\mathcal{L}_{e}}|L-\{e\}|=1+(q+1)(q+1-1)=q^{2}+q+1$.
By double-counting the set of all pairs $(e,L)$ for which $e\in L\in\mathcal{L}$, we see that $|M|(q+1)=(q+1)|\mathcal{L}|$, since each point is in $q+1$ lines and each line has $q+1$ points. Therefore $|\mathcal{L}|=|M|=q^{2}+q+1$.
It remains to argue that $q\geq 2$. Indeed, if $q=1$, then $|M|=3=r(M)$, so $M$ has a coloop, which contradicts that $M$ is connected. ∎
The next lemma shows that modular matroids enjoy a property sometimes known as roundness.
Lemma 4.5.
If $M$ is a connected modular matroid, then every cocircuit of $M$ is spanning.
Proof.
Let $K$ be a cocircuit of $M$, and let $x\in K$. For each $e\notin K$, the rank-two restriction $M|\operatorname{cl}_{M}(\{e,x\})$ is connected by Lemma 4.3, so there is some $f$ for which $T=\{e,f,x\}$ is a triangle of $M$. Since $T$ is a circuit, we have $|T\cap K|\neq 1$, so $f\in K$. It follows that $x\in\operatorname{cl}_{M}(\{e,f\})\subseteq\operatorname{cl}_{M}(K)$. This holds for all $x\in E(M)-K$, so $E(M)-K\subseteq\operatorname{cl}_{M}(K)$, and the lemma follows. ∎
Our last lemma in the section exploits the abundance of triangles in modular matroids to find such a triangle that hits three particular sets.
Lemma 4.6.
Let $H$ be a hyperplane of a connected modular matroid $M$, and let $B$ be a basis of $H$. If $(X,Y)$ is a partition of E(M)HE(M)-H and the sets $B,X$ and $Y$ are nonempty, then $M$ has a triangle intersecting $X,Y$ and $B$.
Proof.
[…]
Suppose, therefore, that $r(M)\geq 3$, and inductively that the lemma holds for matroids of rank r(M)1r(M)-1. Let A=E(M)HA=E(M)-H, let $e\in A$, and let $f\in B$. Let $H_{0}=\operatorname{cl}_{M}(B-f)$ and let $A_{0}=\operatorname{cl}_{M}(H_{0}\cup\{e\})-H_{0}$, noting that $A_{0}\subseteq A$ and $H_{0}\cup A_{0}$ is a hyperplane of $M$. If $X\cap A_{0}$ and $Y\cap A_{0}$ are both nonempty, then the inductive hypothesis applied to the basis BfB-f of $H_{0}$ in $M|(H_{0}\cup A_{0})$ with the partition $(X\cap A_{0},Y\cap A_{0})$ gives the required triangle. We may therefore assume, say, that $X\cap A_{0}=\varnothing$, so $A_{0}\subseteq Y$.
Let $x\in X$, so $x\in A-A_{0}$. By modularity, the line $\operatorname{cl}_{M}(\{x,f\})$ intersects the hyperplane $A_{0}\cup H_{0}$ in a nonloop $y$; since $f\in H-H_{0}$ and $x\in A-A_{0}$, we know that $y\notin\{x,f\}$, so $\{x,y,f\}$ is a triangle. Since $\{x,f\}\cap A=\{x\}$, but the circuit $\{x,y,f\}$ does not intersect the cocircuit $A$ in exactly one element, we have $y\in A$, so $y\in A_{0}\subseteq Y$. This implies that $\{x,y,f\}$ is the required triangle. ∎
5. Coloured Matroids and Decomposers
A coloured matroid is a pair $(M;c)$, where $M$ is a matroid and $c$ is a colouring of $E(M)$. When discussing matroid properties such as connectedness, modularity, or independence of a set, we identify $(M;c)$ with the matroid $M$. The colouring $c$ is simple (with respect to $M$) if $M$ has no loop, and $c(e)=c(f)$ for all $e,f$ that are parallel in $M$. Note that $M$ itself need not be simple for this to hold, but simplicity of $M$ implies simplicity of $c$. A set $X$ is $c$-monochromatic if $c$ is constant on $X$, is $c$-polychromatic otherwise, and is $c$-rainbow if $c$ is injective on $X$. We omit the reference to $c$ if it is clear from context.
The following lemma shows that the existence of a rainbow triangle is equivalent to the existence of a larger rainbow circuit.
Lemma 5.1.
A modular coloured matroid $(M;c)$ has a rainbow triangle if and only if $(M;c)$ has a rainbow circuit of size at least $3$.
Proof.
One direction is immediate. For the other direction, let $C$ be a minimal rainbow circuit of $(M;c)$ for which $|C|\geq 3$. Towards a contradiction, suppose that $|C|\geq 4$.
Let $(X_{1},X_{2})$ be a partition of $C$ with $|X_{1}|,|X_{2}|\geq 2$. Since each $X_{i}$ is a proper subset of $C$, it is independent in $M$, which implies that $r_{M}(\operatorname{cl}_{M}(X_{1})\cap\operatorname{cl}_{M}(X_{2}))=\sqcap_{M}(X_{1},X_{2})=|X_{1}|+|X_{2}|-r_{M}(C)=1$,
Let $e$ be a nonloop in $\operatorname{cl}_{M}(X_{1})\cap\operatorname{cl}_{M}(X_{2})$. For each $i$, the set $X_{i}\cup\{e\}$ is dependent, so contains a circuit $C_{i}$ with $e\in C_{i}$. Note that $|C_{i}|\geq 3$ for each $i$; otherwise, applying circuit elimination to $C_{1}$ and $C_{2}$ yields a circuit properly contained in $C$ (since $|X_{1}|,|X_{2}|\geq 2$). Now, since $C_{i}-\{e\}\subseteq X_{i}$ and $X_{1}\cup X_{2}$ is rainbow, one of the circuits $C_{1},C_{2}$ must be rainbow. But both are clearly smaller than $C$, which contradicts minimality. ∎
We now present several equivalent characterizations of modular coloured matroids without rainbow triangles.
Theorem 5.2.
Let $(M;c)$ be a modular coloured matroid for which $c$ is simple. The following are equivalent:
  1. (i)
    $(M;c)$ has no rainbow triangle;
  2. (ii)
    $(M;c)$ has no rainbow circuit;
  3. (iii)
    $|c(X)|\leq r_{M}(X)$ for all $X\subseteq E(M)$;
  4. (iv)
    $|c(F)|\leq r_{M}(F)$ for every flat $F$ of $M$.
Proof.
By the simplicity of $c$, every rainbow circuit of $M$ has size at least $3$, so the equivalence of (i) and (ii) follows from Lemma 5.1. It is clear that (iii) implies (iv). Moreover, if (iv) holds and $X\subseteq E(M)$, then $|c(X)|\leq|c(\operatorname{cl}_{M}(X))|\leq r_{M}(\operatorname{cl}_{M}(X))=r_{M}(X)$, so (iv) implies (iii). Since a rainbow triangle $T$ satisfies $|c(T)|>r_{M}(T)$, we also know that (iii) implies (i). It is thus sufficient to show that (ii) implies (iii).
Suppose that (ii) holds and (iii) does not, and let $X$ be a minimal set for which $|c(X)|>r_{M}(X)$. For each $e\in X$, the minimality of $X$ implies that
$|c(X-e)|\leq r_{M}(X-e)\leq r_{M}(X)\leq|c(X)|-1\leq|c(X-e)|,$
so equality holds throughout, meaning that $r_{M}(X)=|c(X)|-1$ and that $e$ is the only element of $X$ with colour $c(e)$. This holds for all $e\in X$, so $X$ is rainbow, and $r_{M}(X)=|c(X)|-1=|X|-1$. Therefore $X$ contains a rainbow circuit $C$, contradicting (ii). ∎
A rainbow-triangle-free coloured matroid is essentially the same as one where every line contains at most two colours. A natural generalization is the class of matroids where all rank-$k$ flats contain at most $k$ colours for some fixed $k$. However, we can see by way of the following easy result that this does not make the classification problem any more interesting.
Theorem 5.3.
Let $(M;c)$ be a coloured matroid, and $k\geq 2$ be an integer. If $r(M)\geq k$ and $|c(F)|\leq k$ for every rank-$k$ flat $F$ of $M$, then either $|c|\leq k$, or $(M;c)$ has no rainbow triangle.
Proof.
Suppose for a contradiction that $(M;c)$ has a rainbow triangle $T$, but also that $|c|>k$. Then there is a (k2)(k-2)-element set $S\subseteq E(M)-T$ such that $S\cup T$ is rainbow. Since
$r_{M}(S\cup T)\leq r_{M}(T)+|S|=2+(k-2)=k,$
$S\cup T$ is contained in a rank-$k$ flat of $M$ using at least $|S\cup T|=k+1$ colours, a contradiction. ∎
Our main structural result, Theorem 6.8, adds a nontrivial fifth condition to the list in Theorem 5.2. To be able to state and prove this condition, we need a notion of decomposition for coloured matroids. A decomposer of a coloured matroid $(M;c)$ is a proper flat $F$ of $M$ such that, for all $e\in E(M)-F$, the set $\operatorname{cl}_{M}(F\cup\{e\})-F$ is $c$-monochromatic.
It is easy to see that, if $M$ has no loop, the flat $\varnothing$ is a decomposer if and only if $c$ is simple. A decomposer of rank at least one is nontrivial. For nearly all of our theorems, the matroids addressed are connected and have rank at least one. This means they are loopless, that points are the same as parallel classes, and that nontrivial decomposers are precisely nonempty decomposers.
However, our inductive strategy involves contractions that create parallel pairs; to facilitate this, we work at a level of generality that includes non-simple matroids. The reader should bear in mind the possibility that two distinct elements are parallel (and therefore not in a rainbow triangle), which can affect the course of our arguments.
For a colouring $c$ of a set $X$, we write $c|Y$ for the restriction of $c$ to domain $Y\subseteq X$. If $X\subseteq E(M)$, then we write $(M;c)/X$ for the coloured matroid $((M/X);c|E(M/X))$, and $M|X$ for the coloured matroid $(M|X;c|X)$.
Lemma 5.4.
Let $(M;c)$ be a coloured matroid. If $F$ is a decomposer of $(M;c)$, and $F_{0}$ is a decomposer of $(M;c)|F$, then $F_{0}$ is also a decomposer of $(M;c)$.
Proof.
Clearly $F_{0}$ is a proper flat of $M$. Let $e\in E(M)-F_{0}$. If $e\in F$, then $\operatorname{cl}_{M}(F_{0}\cup\{e\})-F_{0}=\operatorname{cl}_{M|F}(F_{0}\cup\{e\})-F_{0}$, which is monochromatic. If $e\notin F$, then $\operatorname{cl}_{M}(F_{0}\cup\{e\})\cap F=F_{0}$, and so $\operatorname{cl}_{M}(F_{0}\cup\{e\})-F_{0}\subseteq\operatorname{cl}_{M}(F\cup\{e\})-F$, which is monochromatic, as required. ∎
Lemma 5.5.
Let $F$ be a flat of a coloured matroid $(M;c)$. If $F_{0}$ is a decomposer of $(M;c)/F$, then $\operatorname{cl}_{M}(F\cup F_{0})$ is a decomposer of $(M;c)$.
Proof.
Since $F_{0}$ is a proper flat of $M/F$, we see that $\operatorname{cl}_{M}(F\cup F_{0})$ is a proper flat of $M$. Now let $e\in E(M)-\operatorname{cl}_{M}(F\cup F_{0})$. Then
$\operatorname{cl}_{M}(F\cup F_{0}\cup\{e\})-\operatorname{cl}_{M}(F\cup F_{0})\subseteq\operatorname{cl}_{M/F}(F_{0}\cup\{e\})-F_{0},$
which is monochromatic. ∎
We now proceed to show that every connected, modular rainbow-triangle-free coloured matroid of rank at least two has a decomposer. We first handle the rank-two case, which is an easy case analysis.
Lemma 5.6.
Every connected coloured matroid $(M;c)$ with $r(M)=2$ and $|c|\geq 3$ has a rainbow triangle or a nonempty decomposer.
Proof.
Suppose that neither holds. Let $\mathcal{P}$ be the set of points of $M$, and for each $i$, let $\mathcal{P}_{i}=\{P\in\mathcal{P}:|c(P)|=i\}.$ If there exist $P,Q\in\mathcal{P}_{1}$ with $c(P)\neq c(Q)$, then since $|c|\geq 3$ there is some $x$ with $c(x)\notin c(P)\cup c(Q)$, and $\{e,f,x\}$ is a rainbow triangle for any $e\in P$ and $f\in Q$, a contradiction. So any two points in $\mathcal{P}_{1}$ have the same colour.
It follows that, if $\mathcal{P}-\{P\}\subseteq\mathcal{P}_{1}$ for some $P\in\mathcal{P}$, then $P$ is a decomposer of $M$, a contradiction. So there are distinct points $P,Q\in\mathcal{P}\backslash\mathcal{P}_{1}$. Note that for all $(\kappa,\lambda)\in c(P)\times c(Q)$ with $\kappa\neq\lambda$, we have $c(E-P\cup Q)\subseteq\{\kappa,\lambda\}$, as otherwise there is a rainbow triangle. Since $P,Q\notin\mathcal{P}_{q}$, such a pair $(\kappa,\lambda)$ exists, so $c]\subseteq c(P\cup Q)$ and therefore $|c(P\cup Q)|\geq 3$. It follows in turn that $c(P)\times c(Q)$ contains distinct pairs together comprising at least three distinct colours $\kappa,\lambda,\mu$, which implies that $c(E-P\cup Q)\subseteq\{\kappa,\lambda\}\cap\{\kappa,\mu\}\cap\{\lambda,\mu\}=\varnothing$, so $E-P\cup Q=\varnothing$, which contradicts the connectedness of $M$. ∎
We now prove a lemma to facilitate induction in the rank-three case.
Lemma 5.7.
Let $(M;c)$ be a connected, modular coloured matroid with no rainbow triangle. If $r(M)=3$ and $|c|\geq 3$, then $M$ has a line $L$ for which $(M;c)|L$ has a nonempty decomposer.
Proof.
Suppose not, and let $M$ be a counterexample for which $|M|$ is minimized. Let $\mathcal{L}$ be the set of lines of $M$. By Lemma 5.6, we have $|c(L)|\leq 2$ for all $L\in\mathcal{L}$.
5.7.1.
The colouring $c$ is simple.
Subproof:.
If some point $P$ is polychromatic, then since $|c|\geq 3$, there is some $x$ for which $|c(P\cup\{x\})|\geq 3$, and now any line $L$ containing $P$ and $x$ satisfies $|c(L)|\geq 3$, a contradiction. ∎
If $\{e,f\}$ is a parallel pair of $M$, then since $c(e)=c(f)$, the matroid $\e(M;c)\backslash$ e is also a counterexample, a contradiction to minimality. Therefore $M$ is simple. Thus, Lemma 4.4 gives that there is some $q\geq 2$ such that every line of $M$ contains $q+1$ elements, each element of $M$ lies in $q+1$ lines, and $|E(M)|=|\mathcal{L}|=q^{2}+q+1$.
5.7.2.
For each $L\in\mathcal{L}$, there are distinct colours $i,j$ such that $c(L)=\{i,j\}$, and $L$ contains at least two elements of each colour.
Subproof:.
We know that $|c(L)|\leq 2$, but since $(M;c)|L$ has no rank-$1$ decomposer, we also have $|c(L-e)|\geq 2$ for each $e\in L$. The claim follows. ∎
5.7.3.
$|c|=3$.
Subproof:.
If $|c|\geq 4$, we can find $L,L^{\prime}\in\mathcal{L}$ with $|c(L\cup L^{\prime})|\geq 4$. But $|c(L\cup L^{\prime})|=|c(L)|+|c(L^{\prime})|-|c(L\cap L^{\prime})|\leq 2+2-1=3$. ∎
Assume for simplicity that $c(M)=[3]$. For each $i\in[3]$, let $X_{i}=c^{-1}(\{i\})$, so $|X_{1}|+|X_{2}|+|X_{3}|=|M|=q^{2}+q+1$.
For all distinct $i,j\in[3]$, write $\mathcal{L}_{ij}$ for the set of lines $L\in\mathcal{L}$ with $c(L)=\{i,j\}$. By 5.7.2, we have $|\mathcal{L}_{12}|+|\mathcal{L}_{23}|+|\mathcal{L}_{13}|=|\mathcal{L}|=q^{2}+q+1$. If $L_{ij}$ is empty for some $i\neq j$, then either $X_{i}=\varnothing$ or $X_{j}=\varnothing$, contradicting $|c|=3$. Therefore $|\mathcal{L}_{ij}|>0$ for all $i\neq j$.
5.7.4.
Let $i,j,j^{\prime}\in[3]$ be distinct. If $L\in\mathcal{L}_{ij}$ and $L^{\prime}\in\mathcal{L}_{ij^{\prime}}$, then $|L\cap X_{i}|=|L^{\prime}\cap X_{i}|$.
Subproof:.
By symmetry, it suffices to show that $|L\cap X_{i}|\leq|L^{\prime}\cap X_{i}|$. Let $\{e\}=L\cap L^{\prime}$ (so $c(e)=i$), and let $f\in L$ with $c(f)=j$ and $f^{\prime}\in L^{\prime}$ with $c(f^{\prime})=j^{\prime}$. Then $\operatorname{cl}_{M}(\{f,f^{\prime}\})\in\mathcal{L}_{jj^{\prime}}$. By 5.7.2, there exists $g\in\operatorname{cl}_{M}(\{f,f^{\prime}\})-\{f,f^{\prime}\}$ with $c(g)=j$. Note that $g\notin L\cup L^{\prime}$.
Define a function $\varphi:L\to L^{\prime}$ by setting $\varphi(x)$ to be the unique element in $L^{\prime}\cap\operatorname{cl}_{M}(\{x,g\})$. (The set $\operatorname{cl}_{M}(\{x,g\})$ is always a line distinct from $L^{\prime}$, since $g\notin L\cup L^{\prime}$.) If $\varphi(x)=\varphi(x^{\prime})=y$, then the line $\operatorname{cl}_{M}(\{g,y\})$ intersects $L$ in both $x$ and $x^{\prime}$, which implies that $x=x^{\prime}$. Thus $\varphi$ is injective.
Note that $\varphi(e)=e$. If $x\in L-e$ with $c(x)=i$, then since $\{x,g,\varphi(x)\}$ is a triangle and $c(g)=j$, we have $c(\varphi(x))\in\{i,j\}$, and since $c(\varphi(x))\in L^{\prime}\in\mathcal{L}_{ij^{\prime}}$, it follows that $c(\varphi(x))=i$.
Thus, $\varphi(L\cap X_{1})\subseteq(L^{\prime}\cap X_{1})$, and the claim follows by injectivity. ∎
5.7.5.
If $i\neq j$ and $L,L^{\prime}\in\mathcal{L}_{ij}$, then $|L\cap X_{i}|=|L^{\prime}\cap X_{i}|$.
Proof.
Let $k\in[3]-\{i,j\}$, and let $L_{0}\in\mathcal{L}_{ik}$. By 5.7.4, we have $|L\cap X_{i}|=|L_{0}\cap X_{i}|=|L^{\prime}\cap X_{i}|$, as required. ∎
By the last two claims, there exist integers $n_{1},n_{2},n_{3}$ so that, for all $i\neq j$ and $L\in\mathcal{L}_{ij}$, we have $|L\cap X_{i}|=n_{i}$ and $|L\cap X_{j}|=n_{j}$. It follows that $n_{1}+n_{2}=n_{1}+n_{3}=n_{2}+n_{3}=q+1$, whence $n_{1}=n_{2}=n_{3}=\tfrac{1}{2}(q+1)$. So every line in $\mathcal{L}_{ij}$ contains exactly $\tfrac{1}{2}(q+1)$ points of each colour.
Let $t\in[3]$ be such that $|X_{t}|$ is minimized, so $|X_{t}|\leq\tfrac{1}{3}(q^{2}+q+1)$. Let $x\in X_{t}$. Each of the $q+1$ lines through $x$ is in $\mathcal{L}_{ti}$ for some $i\neq t$, so contains $\tfrac{q+1}{2}-1$ points in $X_{t}$ other than $x$. These lines intersect only at $x$, so
$\tfrac{1}{3}(q^{2}+q+1)\geq|X_{t}|\geq 1+(q+1)\left(\tfrac{q+1}{2}-1\right)=\tfrac{1}{2}(q^{2}+1),$
which yields the contradiction $q\leq 1$. ∎
We now show that a decomposer for a hyperplane can be contracted while preserving the property of having no rainbow triangle.
Lemma 5.8.
Let $(M;c)$ be a connected, modular coloured matroid with no rainbow triangle. Let $H$ be a hyperplane of $M$. If $F$ is a decomposer of $(M;c)|H$, then $(M;c)/F$ has no rainbow triangle.
Proof.
By Lemma 4.3, the matroid $M/F$ is connected and modular. Suppose that $(M;c)/F$ has a rainbow triangle $T=\{x,y,z\}$, chosen so that $z\in H$ if possible. Since HFH-F is a hyperplane of the modular matroid $M/F$, there exists $w$ in $\operatorname{cl}_{M/F}(T)\cap H$, and since $T$ is rainbow, the set $T\cup\{w\}$ contains a rainbow triangle of $(M;c)/F$ containing $w$. Thus there exists a rainbow triangle of $M/F$ intersecting $H$; by the choice of $T$, it follows that $z\in H$. Now
$\sqcap_{M}(\{x,y\},F\cup\{z\})$ $=r_{M}(\{x,y\})-r_{M/(F\cup\{z\})}(\{x,y\})$
$=2-(r_{M/F}(T)-1)=1,$
so by modularity, there exists $z^{\prime}\in\operatorname{cl}_{M}(F\cup\{z\})\cap\operatorname{cl}_{M}(\{x,y\})$. If we had $z^{\prime}\in F$, then since $z^{\prime}\in\operatorname{cl}_{M}(\{x,y\})$, we would have
$2=r_{M}(\{x,y\})>r_{M/z^{\prime}}(\{x,y\})\geq r_{M/F}(\{x,y\})=2,$
a contradiction. Therefore $z^{\prime}\in\operatorname{cl}_{M}(F\cup\{z\})-F$, and so $z$ and $z^{\prime}$ are parallel in $M/F$, implying that $z^{\prime}$ is parallel to neither $x$ nor $y$ in $M$. Since $z^{\prime}\in\operatorname{cl}_{M}(\{x,y\})$, It follows that $\{x,y,z^{\prime}\}$ is a triangle of $M$.
Because $F\cup\{z\}\subseteq H$ and $F$ is a decomposer of $(M;c)|H$, the set $\operatorname{cl}_{M}(F\cup\{z\})-F$, which contains $z$ and $z^{\prime}$, is monochromatic. Therefore $c(z^{\prime})=c(z)$. Since $\{x,y,z\}$ is rainbow, so is the triangle $\{x,y,z^{\prime}\}$ of $M$, giving a contradiction. ∎
We now prove the decomposition theorem that will imply our main structure theorem for rainbow-triangle-free matroids.
Theorem 5.9.
Let $(M;c)$ be a connected, modular coloured matroid with no rainbow triangle. If $r(M)\geq 2$ and $|c|\geq 3$, then $(M;c)$ has a nonempty decomposer.
Proof.
Suppose not, and let $(M;c)$ be a counterexample chosen so that $r(M)$ is minimized.
5.9.1.
There is a hyperplane $H$ of $M$ for which $(M;c)|H$ has a nonempty decomposer.
Subproof:.
If $M$ has a hyperplane $H$ for which $|c(H)|\geq 3$, then by the minimality in the choice of $M$, the restriction $(M;c)|H$ has a proper nonempty decomposer, as required.
Assume, therefore, that every hyperplane $H$ satisfies $|c(H)|\leq 2$. Let $x_{1},x_{2},x_{3}$ be elements of different colours. Since no hyperplane contains all three, we have $r(M)\leq r(\{x_{1},x_{2},x_{3}\})\leq 3$. If $r(M)=2$, then Lemma 5.6 contradicts the choice of $M$ as a counterexample. So $r(M)=3$, and the claim follows from Lemma 5.7. ∎
Let $H$ be a hyperplane of $M$ for which $(M;c)|H$ has a nonempty decomposer, and let $F$ be a minimal nonempty decomposer of $(M;c)|H$.
5.9.2.
$(M;c)|F$ has no nonempty decomposer.
Subproof:.
If $F_{0}$ were such a decomposer, then since $F$ is a decomposer of $(M;c)|H$, by Lemma 5.4 we see that $F_{0}$ is a decomposer of $(M;c)|H$ that is properly contained in $F$, a contradiction to the minimality in the choice of $F$. ∎
Since $r(M|F)<r(M)$, and the modular, connected coloured matroid $(M;c)|F$ has no rainbow triangle but is not a counterexample, the minimality in the choice of $M$ immediately gives
5.9.3.
$r_{M}(F)=1$ or $|c(F)|\leq 2$.
Since $F$ is not a decomposer of $M$, there exists $e\in E(M)-F$ so that the set $A=\operatorname{cl}_{M}(F\cup\{e\})-F$ contains elements of two different colours. Note that $F\cup A$ is a flat and that $F$ is a hyperplane of $M|F\cup A$.
5.9.4.
$c(F)\subseteq c(A)$.
Subproof:.
Let $\beta\in c(F)$. If $c^{-1}(\beta)$ is not spanning in $F$, then $r_{M}(F)\geq 2$ and $F$ has a hyperplane $F_{0}$ such that $c(F-F_{0})\subseteq c(F)-\{\beta\}$, which by 5.9.3 implies that $F-F_{0}$ is monochromatic. This implies that $F_{0}$ is a nonempty decomposer of $(M;c)|F$, a contradiction. Therefore $c^{-1}(\beta)$ is spanning in $F$. Let $B$ be a basis for $F$ with $c(B)=\{\beta\}$.
By its choice, the set $A$ is not monochromatic. Let $(X,Y)$ be a partition of $A$ with $X,Y$ nonempty so that $c(X)$ and $c(Y)$ are disjoint. By Lemma 4.6, there is a triangle of $M$ intersecting $X,Y$ and $B$. This triangle is not rainbow, which implies that $\beta\in c(X)\cup c(Y)\subseteq c(A)$. This holds for all $\beta\in c(F)$, so the claim follows. ∎
5.9.5.
$(M;c)/F$ has a nonempty decomposer.
Subproof:.
By Lemma 5.8, we see that $(M;c)/F$ has no rainbow triangle. Since $F$ is a proper subflat of the hyperplane $H$, we have $r(M/F)\geq 2$, so by induction it suffices to show that $c(M/F)=c(M)$. Indeed, since $A\subseteq E(M)-F$, we have $c(M/F)\supseteq c(A)\supseteq c(F)$ by 5.9.4, so $c(M/F)=c(M/F)\cup c(F)=c(M)$. ∎
Now Lemma 5.5 implies that $M$ has a nonempty decomposer, contrary to its choice as a counterexample. ∎
6. Lift-Joins
To turn Theorem 5.9 into a structure theorem, we need an operation that can be used to glue two rainbow-triangle-free colourings together into a larger one.
We now describe this operation, then use Theorem 5.9 to derive Theorem 6.7. Theorem 5.9 is the hard part of the actual proof; the work in this section amounts to straightforward manipulation of definitions in order that Theorem 6.7 can be carefully stated and proved.
We call the gluing operation a lift-join. This was first defined in a slightly different (two-coloured) context in [6]. The definition uses the fact that, if $F_{1}$ and $F_{2}$ are skew flats of a modular matroid $M$ and $e\in\operatorname{cl}_{M}(F_{1}\cup F_{2})-F_{1}$, then the flats $\operatorname{cl}_{M}(F_{1}\cup\{e\})$ and $F_{2}$ intersect in a point. A lift-join combines colourings of $F_{1}$ and $F_{2}$ to obtain a colouring of the closure of their union.
Definition 6.1.
Let $F_{1}$ and $F_{2}$ be skew flats of a modular matroid $M$, and let $c_{1},c_{2}$ be colourings of $F_{1},F_{2}$ respectively with $c_{2}$ simple in $M$.
The lift-join $c_{1}\otimes c_{2}$ is the colouring $c$ of $\operatorname{cl}_{M}(F_{1}\cup F_{2})$ such that $c(e)=c_{1}(e)$ for all $e\in F_{1}$, and for each $e\in\operatorname{cl}_{M}(F_{1}\cup F_{2})-F_{1}$, we have $\{c(e)\}=c_{2}(\operatorname{cl}_{M}(F_{1}\cup\{e\})\cap F_{2})$.
We establish some basic properties of lift-joins.
Lemma 6.2.
Let $F_{1}$ and $F_{2}$ be skew flats of a modular matroid $M$, and let $c_{1},c_{2}$ be colourings of $F_{1},F_{2}$ respectively with $c_{2}$ simple. Then
  1. (i)
    $c_{1}\otimes c_{2}]=[c_{1}]\cup[c_{2}]$,
  2. (ii)
    If $c_{1}$ is simple in $M$, then so is $c_{1}\otimes c_{2}$,
  3. (iii)
    $(c_{1}\otimes c_{2})|F_{i}=c_{i}$ for each $i\in[2]$,
  4. (iv)
    $d\circ(c_{1}\otimes c_{2})=(d\circ c_{1})\otimes(d\circ c_{2})$ for every function $d$ with domain containing $c_{1}]\cup[c_{2}]$,
  5. (v)
    If $r_{M}(F_{2})>0$, then $F_{1}$ is a decomposer of the coloured matroid $(M|\operatorname{cl}_{M}(F_{1}\cup F_{2});c_{1}\otimes c_{2})$.
Proof.
The first four facts follow immediately from the definition.
To see (v), suppose that $r(F_{2})>0$, so $F_{1}$ is a proper flat of $M|\operatorname{cl}_{M}(F_{1}\cup F_{2})$. Let $x\in\operatorname{cl}_{M}(F_{1}\cup F_{2})-F_{1}$, and $y$ be a nonloop in $\operatorname{cl}_{M}(F_{1}\cup\{x\})\cap F_{2}$. By construction, we have $c(x^{\prime})=c(y)$ for all $x^{\prime}\in\operatorname{cl}_{M}(F_{1}\cup\{x\})-F_{1}$, so $\operatorname{cl}_{M}(F_{1}\cup\{x\})-F_{1}$ is monochromatic, as required. ∎
In fact, appropriately phrased, the implication in Lemma (6.2)(v) holds in both directions.
Lemma 6.3.
Let $(M;c)$ be a modular coloured matroid, and $F_{1}$ and $F_{2}$ be complementary flats of $M$. The following are equivalent:
  • $F_{1}$ is a decomposer of $(M;c)$,
  • $F_{1}$ is a proper flat, $c|F_{2}$ is simple, and $c=(c|F_{1})\otimes(c|F_{2})$.
Proof.
Suppose first that $F_{1}$ is a decomposer of $(M;c)$. By definition, $F_{1}$ is a proper flat. If $e,e^{\prime}\in F_{2}$ are parallel, then since $e^{\prime}\in\operatorname{cl}_{M}(F_{1}\cup\{e\})-F_{1}$, we have $c(e)=c(e^{\prime})$, so $c|F_{2}$ is simple. For each $e\in F_{1}$ we have $c(e)=(c|F_{1})(e)$. For each $e\in\operatorname{cl}_{M}(F_{1}\cup F_{2})-F_{1}$, let $P=\operatorname{cl}_{M}(F_{1}\cup\{e\})\cap F_{2}$; this is a point of $M$. Since $P\subseteq\operatorname{cl}_{M}(F_{1}\cup\{e\})$, we know that $(c|F_{2})(P)=c(P)=\{c(e)\}$, and therefore $c=(c|F_{1})\otimes(c|F_{2})$.
Conversely, suppose that $F_{1}$ is a proper flat, $c|F_{2}$ is simple, and $c=(c|F_{1})\otimes(c|F_{2})$. Since $r_{M}(F_{2})=r(M)-r_{M}(F_{1})>0$ and $\operatorname{cl}_{M}(F_{1}\cup F_{2})=E(M)$, Lemma 6.2(v) implies that $F_{1}$ is a decomposer of $M$. ∎
Recall that sets $X_{1},\dotsc,X_{t}$ are mutually skew in $M$ if $r_{M}(\cup_{i}X_{i})=\sum_{i}r_{M}(X_{i})$, or equivalently if $X_{i}$ is skew in $M$ to $\cup_{j<i}X_{j}$ for all $i\in\{1,\dotsc,t\}$. If $c_{1},\dotsc,c_{t}$ are colourings of mutually skew flats $F_{1},\dotsc,F_{t}$ of $M$ such that $c_{i}$ is simple for all $i>1$, then we can iterate lift-joins to obtain a colouring $\bigotimes_{i}c_{i}$ of $\operatorname{cl}_{M}(\cup_{i}F_{i})$. More precisely, we have $\bigotimes_{i}c_{i}=(\dotsc(((c_{1}\otimes c_{2})\otimes c_{3})\otimes\dotsc)\otimes c_{t})$. However, the next lemma proves associativity, telling us that the parentheses can be omitted. Note that the right-hand-side is well-defined by Lemma 6.2(ii).
Lemma 6.4.
Let $F_{1},F_{2},F_{3}$ be mutually skew flats of a modular matroid $M$, and $c_{1},c_{2},c_{3}$ be colourings of $F_{1},F_{2},F_{3}$ respectively with $c_{2}$ and $c_{3}$ simple. Then $(c_{1}\otimes c_{2})\otimes c_{3}=c_{1}\otimes(c_{2}\otimes c_{3})$.
Proof.
Let $e\in\operatorname{cl}_{M}(F_{1}\cup F_{2}\cup F_{3})$. If $e\in F_{1}$, then $e\in\operatorname{cl}_{M}(F_{1}\cup F_{2})$, so
$((c_{1}\otimes c_{2})\otimes c_{3})(e)=(c_{1}\otimes c_{2})(e)=c_{1}(e)=(c_{1}\otimes(c_{2}\otimes c_{3}))(e).$
If $e\in\operatorname{cl}_{M}(F_{1}\cup F_{2})-F_{1}$, then let $f\in F_{2}$ be parallel to $e$ in $M/F_{1}$. Since $f\in\operatorname{cl}_{M}(F_{2}\cup F_{3})$, we have
$((c_{1}\otimes c_{2})\otimes c_{3})(e)=(c_{1}\otimes c_{2})(e)=c_{2}(f)=(c_{1}\otimes(c_{2}\otimes c_{3}))(e).$
Finally, suppose that $e\in\operatorname{cl}_{M}(F_{1}\cup F_{2}\cup F_{3})-\operatorname{cl}_{M}(F_{1}\cup F_{2})$. Let $f^{\prime}\in\operatorname{cl}_{M}(F_{2}\cup F_{3})$ be parallel to $e$ in $M/F_{1}$. If we have $f^{\prime}\in\operatorname{cl}_{M}(F_{2})$ we would have $e\in\operatorname{cl}_{M/F_{1}}(F_{2})=\operatorname{cl}_{M}(F_{1}\cup F_{2})$, a contradiction. So $f^{\prime}\in\operatorname{cl}_{M}(F_{2}\cup F_{3})-\operatorname{cl}_{M}(F_{2})$. Let $f\in F_{3}$ be parallel to $f^{\prime}$ in $M/F_{2}$.
By definition, we have $(c_{1}\otimes(c_{2}\otimes c_{3}))(f^{\prime})=(c_{2}\otimes c_{3})(f^{\prime})=c_{3}(f)$. On the other hand, since $e$ and $f$ are parallel in $M/F_{1}$, while $f$ and $f^{\prime}$ are parallel in $M/F_{2}$, all three are parallel in $M/\operatorname{cl}_{M}(F_{1}\cup F_{2})$, which gives $((c_{1}\otimes c_{2})\otimes c_{3})(e)=c_{3}(f)=(c_{1}\otimes(c_{2}\otimes c_{3}))(e)$, as required. ∎
It is easy to see, on the other hand, that lift-joins are not in general commutative, so the order of the indices is still important in an expression $\bigotimes_{i}c_{i}$. Note also that for such an expression to be well-defined, each $c_{i}$ except possibly $c_{1}$ must be simple.
Lemma 6.5.
Let $F_{1}$ and $F_{2}$ be skew flats of a modular matroid $M$, and $c_{1},c_{2}$ be colourings of $F_{1},F_{2}$ respectively so that $c_{2}$ is simple. If $(M|\operatorname{cl}_{M}(F_{1}\cup F_{2});c_{1}\otimes c_{2})$ has a rainbow triangle, then so does one of $(M|F_{1};c_{1})$ or $(M|F_{2};c_{2})$.
Proof.
We may assume that $\operatorname{cl}_{M}(F_{1}\cup F_{2})=E(M)$ and that $r_{M}(F_{2})>0$. Suppose that $T$ is a rainbow triangle in $(M;c_{1}\otimes c_{2})$.
If $\sqcap_{M}(F_{1},T)=2$, then $T\subseteq F_{1}$, so $T$ is a rainbow triangle of $(M|F_{1};c_{1})$. If $\sqcap_{M}(T,F_{1})=1$, then let $e$ be a nonloop in $\operatorname{cl}_{M}(T)\cap F_{1}$, and let $x\in T-\operatorname{cl}_{M}(\{e\})$. Since $T\subseteq\operatorname{cl}_{M}(\{e,x\})$ and $T\not\subseteq F_{1}$, we know that $x\notin F_{1}$. Now $T-\operatorname{cl}_{M}(\{e\})\subseteq\operatorname{cl}_{M}(\{e,x\})-F_{1}\subseteq\operatorname{cl}_{M}(F_{1}\cup\{x\})-F_{1}$. But by Lemma 6.2(v), the set $\operatorname{cl}_{M}(F_{1}\cup\{x\})-F_{1}$ is therefore monochromatic and contains at least two elements of $T$, a contradiction.
Suppose, therefore, that $\sqcap_{M}(F_{1},T)=0$, so $T$ is skew to $F_{1}$, and $(M/F_{1})|T=M|T$. Each $e\in T$ is parallel to some element of $F_{2}$ in $M/F_{1}$. Let $T^{\prime}\subseteq T$ be a triple containing an element parallel in $M/F_{1}$ to each element of $T$. Since $F_{1}$ and $F_{2}$ are skew, we have $(M/F_{1})|F_{2}=M|F_{2}$. By the choice of $T^{\prime}$ we have
$M|T^{\prime}=(M/F_{1})|T^{\prime}\cong(M/F_{1})|T=M|T,$
and thus $T^{\prime}$ is a triangle of $M$. But also by the choice of $T^{\prime}$ and Lemma 6.2(v), we see that $(c_{1}\otimes c_{2})(T)=c_{2}(T^{\prime})$, and so $T^{\prime}$ is a rainbow triangle of $(M|F_{2};c_{2})$. ∎
Given a set $X$ of elements of a matroid $M$, write $\omega_{M}(X)$ for the rank of the largest flat of $M$ that is contained in $X$. (In the case where $M\cong\operatorname{PG}(n-1,2)$, this essentially agrees with the definition of $\omega$ given earlier for subsets of $\mathbb{F}_{2}^{n}$.) The following lemma gives an expression for $\omega_{M}(X)$, where $X$ is a colour class in a lift-join of colourings.
Lemma 6.6.
Let $F_{1},\dotsc,F_{t}$ be mutually skew flats of a connected modular coloured matroid $(M;c)$, such that $c=\bigotimes_{i=1}^{t}(c|F_{i})$. If $X$ is a colour class of $c$, then $\omega_{M}(X)=\sum_{i=1}^{t}\omega_{M}(F_{i}\cap X)$.
Proof.
By induction and associativity, it suffices to show this for $t=2$. Suppose, therefore, that $c=(c|F_{1})\otimes(c|F_{2})$ for complementary flats $F_{1},F_{2}$ of $M$, and let $X=c^{-1}(\kappa)$ for some $\kappa\in[c]$.
For each $i$, let $K_{i}\subseteq F_{i}\cap X$ and $r_{M}(K_{i})=\omega_{M}(F_{1}\cap X)$. We clearly have $c(K_{1})=\{\kappa\},$ and for each $e\in\operatorname{cl}_{M}(K_{1}\cup K_{2})-K_{1}$, the point $f\in K_{2}\cap\operatorname{cl}_{M}(K\cup\{e\})$ has $c(f)=\kappa$ and $c(f)=c(e)$, so we have $c(\operatorname{cl}_{M}(K_{1}\cup K_{2}))=\{\kappa\}$ and thus $\omega_{M}(X)\geq r_{M}(K_{1}\cup K_{2})=r_{M}(K_{1})+r_{M}(K_{2})=\sum_{i=1}^{2}\omega_{M}(F_{i}\cap X)$.
For the other direction, let $K$ be a flat of $M$ with $K\subseteq X$ and $r_{M}(K)=\omega_{M}(X)$. Let $K_{1}=K\cap F_{1}$, and let $K_{2}$ be a complement of $K_{1}$ in $M|K$. By construction, we have $\omega_{M}(X)=r_{M}(K_{1})+r_{M}(K_{2})$, and $r_{M}(K_{1})\leq\omega_{M}(K\cap F_{1})$, so it suffices to show that $r_{M}(K_{2})\leq\omega_{M}(F_{2}\cap X)$.
Since $K_{2}$ and $F_{1}$ are disjoint, modularity implies that they are skew in $M$, so $(M/F_{1})|K_{2}=M|K_{2}$. By modularity, for each $e\in K_{2}$, the flat $\operatorname{cl}_{M}(F_{1}\cup\{e\})$ intersects $F_{2}$ in a unique element $\varphi(e)$, and $e$ and $\varphi(e)$ are parallel in $M/F_{1}$. Let $K_{2}^{\prime}=\varphi(K_{2})\subseteq F_{2}$. Since $K_{2}^{\prime}\subseteq F_{2}$ and $F_{2}$ is skew to $F_{1}$, we have $M|K_{2}=(M/F_{1})|K_{2}\cong(M/F_{1})|K_{2}^{\prime}=M|K_{2}^{\prime}$, where the isomorphism is given by $\varphi$. So $K_{2}^{\prime}$ is a flat of $M$ of rank equal to $r_{M}(K_{2})$, and since $e\in\operatorname{cl}_{M}(F_{1}\cup\varphi(e))$ for each $e$, we have $c(e)=c(\varphi(e))$ and thus $c(K_{2}^{\prime})=c(K_{2})\subseteq\{\kappa\}$. It follows that $r_{M}(K_{2})=r_{M}(K_{2}^{\prime})\leq\omega_{M}(F_{2}\cap X)$, as required. ∎
Finally, we can state our main structure theorem, which roughly asserts that rainbow-triangle-free coloured matroids are precisely those that can be obtained by gluing simple two-coloured flats together using lift-joins. The only technical exception is the first summand, which can be non-simple, or a rank-one flat that uses more than two colours.
Theorem 6.7.
Let $(M;c)$ be a connected, modular coloured matroid. The following are equivalent:
  1. (i)
    $(M;c)$ is rainbow-triangle-free;
  2. (ii)
    $(M;c)$ has no rainbow circuit of size $3$ or more;
  3. (iii)
    there are mutually skew nonempty flats $F_{i},\dotsc,F_{t}$ of $M$ such that $c=\bigotimes_{i=1}^{t}(c|F_{i})$, the colouring $c|F_{i}$ is simple with $|c(F_{i})|\leq 2$ for all $i\geq 2$, and either $r_{M}(F_{1})=1$ or $|c(F_{1})|\leq 2$.
Proof.
The equivalence of (i) and (ii) follows from Lemma 5.1, so it suffices to show that (i) and (iii) are equivalent.
First suppose that there are flats $F_{i}$ such that (iii) holds. This implies that $c_{i}:=c|F_{i}$ is rainbow-triangle-free for each $i$. It follows by induction and Lemma 6.5 that $(M;c)=(M;\bigotimes_{i=1}^{t}c_{i})$ is rainbow-triangle-free.
Conversely, suppose that $(M;c)$ is rainbow-triangle-free. If $r(M)=0$, then we can take $t=0$. If $r(M)=1$ or $|c|\leq 2$, then it is easy to check that we can take $t=1$ and $F_{1}=E(M)$. We may assume, therefore, that $|c|\geq 3$ and $r(M)\geq 2$, and that (i) implies (iii) for all matroids of smaller rank than $M$.
By Theorem 5.9, there is a nontrivial decomposer $F^{1}$ of $(M;c)$. Let $F^{2}$ be a complement of $F^{1}$ in $M$. Since $F^{1}$ is a decomposer, we have that $r_{M}(F^{2})>0$ and the restriction $c|F^{2}$ is a simple colouring.
Using the fact that $F^{1}$ and $F^{2}$ are proper flats and induction, for each $i\in[2]$, there are flats $F_{1}^{i},\dotsc,F_{t_{i}}^{i}$ satisfying (iii) in the coloured matroid $(M;c)|F^{i}$. For each $i\in[t_{1}]$, let $c^{1}_{i}:=c|F^{1}_{i}$ and for each $j\in[t_{2}]$, let $c^{2}_{j}=c|F^{2}_{j}$. By Lemma 6.3, $c=(c|F^{1})\otimes(c|F^{2})=(\bigotimes_{i=1}^{t_{1}}c^{1}_{i})\otimes(\bigotimes_{j=1}^{t_{2}}c_{j}^{2})$.
We now show that $F_{1}^{1},\dotsc,F_{t_{1}}^{1},F_{1}^{2},\dotsc,F_{t_{2}}^{2}$ is a sequence of flats satisfying (iii) for $M$. This is almost completely straightforward, since $\otimes$ is associative and the flats in the collection $\{F_{i}^{1}\}\cup\{F_{j}^{2}\}$ are mutually skew. All that remains is to show that the colouring $c|F^{2}_{1}$ is simple and that $|c(F^{2}_{1})|\leq 2$. The simplicity follows from the fact that $c|F^{2}_{1}$ is a restriction of the simple colouring $c|F^{2}$. If $|c(F^{2}_{1})|>2$ then since $F^{2}_{1},\dotsc,F^{2}_{t_{2}}$ satisfy (iii), we know that $r_{M}(F^{2}_{1})=1$, but then simplicity gives $|c(F^{2}_{1})|=1$, a contradiction. ∎
Note that any collection of flats satisfying (iii) must have spanning union in $M$, since otherwise $\bigotimes_{i}(c|F_{i})$ would not be defined on all of $M$.
When the colouring $c$ is itself simple, all the $c_{i}$ must be simple, so we can guarantee that $|c_{i}|\leq 2$ for all $i$, giving a less technical version of Theorem 6.7. We remark that all the other equivalent conditions in Theorem 5.2 could also be added to this list.
Theorem 6.8.
Let $(M;c)$ be a connected, modular coloured matroid for which $c$ is simple. The following are equivalent:
  1. (i)
    $(M;c)$ is rainbow-triangle-free;
  2. (ii)
    there are mutually skew nonempty flats $F_{1},\dotsc,F_{t}$ of $M$ such that $c=\bigotimes_{i=1}^{t}(c|F_{i})$, and $|c(F_{i})|\leq 2$ for each $i$.
In particular, this applies whenever $M$ is the matroid of a finite-dimensional Desarguesian projective geometry over a finite field. The statement of Theorem 6.8 is easy to extend meaningfully to geometries over infinite fields, but in fact the extension is false in general, for interesting reasons; see Section 12 for a discussion.
Finally, we formulate the rank-three case of the above as a theorem about projective planes, which are equivalent to simple, modular, connected rank-three matroids. The case of Desarguesian planes was independently proved by Carter and Vogt [7] and by Hales and Straus [14]. The proof, using Theorem 6.8, is straightforward.
Theorem 6.9.
Let $c$ be a colouring of the points of a projective plane $P$. Then every line of $P$ contains points of at most two colours if and only if either
  1. 1.
    $|c|\leq 2$,
  2. 2.
    there are colours $\alpha,\beta_{1},\beta_{2}$ so that $c]=\{\alpha,\beta_{1},\beta_{2}\}$, and either
    1. i.
      there is a point $x$ of $P$ so that $c(x)=\alpha$, and each line $L$ through $x$ satisfies $c(L-\{x\})=\{\beta_{1}\}$ or $c(L-\{x\})=\{\beta_{2}\}$, or
    2. ii.
      there is a line $L$ of $P$ such that $c(P-L)=\{\alpha\}$ and $c(L)=\{\beta_{1},\beta_{2}\}$, and $L$ contains at least two points of each colour.
Kemperman’s theorem
In the case where $M$ is a Desarguesian projective geometry and $|c|=3$, Theorem 6.8 can almost certainly be derived from Kemperman’s theorem on small sumsets in abelian groups. Though the original version of the theorem in [17] does not use this terminology, we phrase our discussion in terms of the shorter proof in [4], which uses the convenient concept of a critical trio, defined to be a triple $(A,B,C)$ of subsets of an additive abelian group $G$ for which $|A|+|B|+|C|>|G|$ and $0\notin A+B+C$.
[…]
Kemperman’s theorem as formulated in [4, Theorem 4.5] gives an exact characterization of all critical trios, in terms of so-called ‘beats’, ‘chords’, ‘continuations’ and a chain of subgroups. We do not state the result here, since its statement requires a plethora of definitions, but the correspondence with Theorem 6.8 is clear – ‘impure beats’ correspond to lift-joins, and the chain of subgroups correspond to the subspaces spanned by unions of initial segments of the $F_{i}$. Deriving the three-coloured case of Theorem 6.8 from Kemperman’s theorem is likely a matter of little more than reconciling distinct sets of terminologies.
With more than three colours, the simplicity of the correspondence breaks down. While one could possibly use an inductive strategy where distinct colours are merged to find a proof of Theorem 6.8 using Kemperman’s theorem, such a proof could easily be longer than ours.
7. Targets
Before applying Theorem 6.8 to derive our main results on the geometric Erdős-Hajnal conjecture, we define a special type of colouring. This is a version of a definition originally given in [21] in the language of induced submatroids, which is essentially the two-coloured case.
Call a coloured matroid $(M;c)$ a target if there are flats $\operatorname{cl}_{M}(\varnothing)=F_{0}\subseteq F_{1}\subseteq\dotsc\subseteq F_{k}=E(M)$ of $M$ such that every nonempty set of the form $F_{0}$ or $F_{i+1}-F_{i}$ is monochromatic.
Given a target $(M;c)$ with $r(M)\geq 1$, it can be convenient to choose the flats $F_{0}\subseteq\dotsc\subseteq F_{k}$ with $k$ minimal. Given such a choice, $F_{i+1}-F_{i-1}$ must contain at least two colours for all $1\leq i<k$, since otherwise we could remove $F_{i}$ from the list. It follows that there is always a choice of the $F_{i}$ so that the containments $F_{i}\subseteq F_{i+1}$ are all strict, and for all $i$, the sets $F_{i}-F_{i-1}$ and $F_{i+1}-F_{i}$ are monochromatic with different colours.
This section will establish some basic properties of targets, as well as prove two theorems that classify targets in terms of excluded subconfigurations. We omit the trivial proofs of the first two lemmas.
Lemma 7.1.
If $(M;c)$ is a target and $F$ is a flat of $M$, then $(M;c)|F$ is a target.
Lemma 7.2.
If $(M;c)$ is a target and $d$ is a function whose domain contains $[c]$, then $(M;d\circ c)$ is a target.
Lemma 7.3.
A connected, modular coloured matroid $(M;c)$ is a target if and only if $c$ is a lift-join of a family of monochromatic flats of $M$.
Proof.
Suppose that $\varnothing=F_{0}\subseteq\dotsc F_{k}=E(M)$ are flats certifying that $(M;c)$ is a target. For each $1\leq i\leq k$, since $F_{i}-F_{i-1}$ is monochromatic, the flat $F_{i-1}$ is a decomposer of $M|F_{i}$, so by Lemma 6.3, there is a complement $K_{i}$ of $F_{i-1}$ in $M|F_{i}$ for which $c|F_{i}=(c|F_{i-1})\otimes(c|K_{i})$. It follows easily by induction on $t$ that $c|F_{t}=\bigotimes_{i=1}^{t}(c|K_{i})$ for all $1\leq t\leq k$. Setting $k=t$ gives that $c=\bigotimes_{i=1}^{k}(c|K_{i})$, as required.
Conversely, suppose that $c=\bigotimes_{i=1}^{k}(c|K_{i})$, where $K_{1},\dotsc,K_{k}$ are mutually skew monochromatic flats of $M$ with spanning union. It is immediate that the collection of flats $F_{i}:0\leq i\leq k$ defined by $F_{i}=\operatorname{cl}\left(\bigcup_{j=1}^{i}K_{j}\right)$ certifies that $(M;c)$ is a target. ∎
The next lemma relates targets to decomposers.
Lemma 7.4.
Let $F$ be a decomposer of a modular coloured matroid $(M;c)$, and $F^{\prime}$ be a complement of $F$ in $M$. If $(M;c)|F$ and $(M;c)|F^{\prime}$ are targets, then $(M;c)$ is a target.
Proof.
By hypothesis and Lemma 7.3, we know that $c|F$ is a lift-join of a family of monochromatic flats of $M|F$, and likewise for $c|F^{\prime}$ and $M|F^{\prime}$. By Lemma 6.3, we have $c=(c|F)\otimes(c|F^{\prime})$. The result now follows from associativity and Lemma 7.3 applied to $(M;c)$. ∎
Lemma 7.5.
A rank-two coloured matroid $(M;c)$ is a target if and only if $M$ has a point $P$ for which $P$ and E(M)PE(M)-P are monochromatic.
Proof.
If $M$ has such a point $P$, the flats $\operatorname{cl}_{M}(\varnothing)\subseteq P\subseteq E(M)$ certify that $(M;c)$ is a target. If $M$ is a target, let $F_{0},\dotsc,F_{k}$ be flats of $M$ such that $1\leq k\leq 2$ and $\operatorname{cl}_{M}(\varnothing)=F_{0}\subset\dotsc\subset F_{k}=E(M)$ while each set $F_{i+1}-F_{i}$ is monochromatic. If $k=1$ then $M$ is monochromatic and any point $P$ satisfies the lemma. If $k=2$, then $F_{1}$ is a point satisfying the lemma. ∎
Lemma 7.6.
If $(M;c)$ is a modular target for which $|c|\geq 2$, then there are distinct colours $\kappa,\lambda$ and a nontrivial decomposer $F$ of $M$ such that E(M)FE(M)-F is monochromatic with colour $\kappa$, and $M|F$ is a target that has a monochromatic basis of colour $\lambda$.
Proof.
Let $k$ be minimal so that there are flats $F_{0},\dotsc,F_{k}$ of $M$ with $\operatorname{cl}_{M}(\varnothing)=F_{0}\subset\dotsc\subset F_{k}=E(M)$ and each $F_{i+1}-F_{i}$ monochromatic. Since $|c|\geq 2$ we have $k\geq 2$. Let $\{\kappa\}=c(E(M)-F_{k-1})$ and $\{\lambda\}=c(F_{k-1}-F_{k-2})$. If $\kappa=\lambda$, we could remove $F_{k-1}$ from the sequence, which contradicts minimality, so $\kappa\neq\lambda$. Since $F_{k-2}$ is a proper subflat of $F_{k-1}$, the set $F_{k-1}-F_{k-2}$ is codependent in $M|F_{k-1}$, so it contains a cocircuit, which by Lemma 4.5 is spanning. It therefore contains a monochromatic basis of $F_{k-1}$ with colour $\lambda$, so $F=F_{k-1}$ satisfies the lemma. ∎
Lemma 7.7.
Let $(M;c)$ be a connected, modular coloured matroid with $r(M)\geq 3$ that is not a target, but for which $(M;c)|H$ is a target for every hyperplane $H$ of $M$. Then
  1. (i)
    $(M;c)|F$ is a target for every proper flat $F$ of $M$,
  2. (ii)
    $M$ has no nontrivial decomposer, and
  3. (iii)
    $|c|=2$.
Proof.
Since every proper subflat is contained in a hyperplane, (i) is immediate by Lemma 7.1.
To see (ii), let $F$ be a nontrivial decomposer of $M$, and $F^{\prime}$ be a complement of $F$ in $M$. We have $r_{M}(F^{\prime})=r(M)-r_{M}(F)<r(M)$, so both $F$ and $F^{\prime}$ are proper subflats and are therefore targets by (i). It follows from Lemma 7.4 that $M$ is a target, a contradiction.
Since $M$ is not a target, we have $|c|\geq 2$. If $M$ has a rainbow triangle $T$, then $(M;c)|\operatorname{cl}_{M}(T)$ is not a target by Lemma 7.5, contrary to (i). So $M$ has no rainbow triangle. If $|c|\geq 3$, it thus follows from Theorem 5.9 that $(M;c)$ has a nontrivial decomposer $F$, contradicting (ii). ∎
This first theorem shows that a modular coloured matroid whose planes are targets is a target. In the special case of two-coloured binary projective geometries, this was essentially proved in [21].
Theorem 7.8.
Let $(M;c)$ be a connected modular coloured matroid with $r(M)\geq 3$. Then $(M;c)$ is a target if and only if $(M;c)|P$ is a target for every plane $P$ of $M$.
Proof.
Let $(M;c)$ be a counterexample of minimum rank, so $(M;c)$ is not a target, but restricting to any plane gives a target, and hence $r(M)\geq 4$. By this minimality, the coloured matroid $(M;c)|H$ is a target for each hyperplane $H$ of $M$, so $(M;c)$ satisfies the hypotheses of Lemma 7.7.
By Lemma 7.7(iii), we have $|c|=2$. Let $H$ be a hyperplane of $M$ for which $|c(H)|=2$, noting that $(M;c)|H$ is a target. By the choice of $H$ and Lemma 7.6, there is a nontrivial decomposer $F$ of $(M;c)|H$ and a basis $B$ of $F$ such that $c(B)=\{\lambda\}$ and $c(H-F)=\{\kappa\}$ for distinct colours $\kappa,\lambda$.
Since $F$ is not a decomposer of $M$, there exists $e\in E(M)-F$ such that the set $A=\operatorname{cl}_{M}(F\cup\{e\})-F$ satisfies $c(A)=\{\kappa,\lambda\}$. Note that $F\cup A$ is a flat, and $F$ is a hyperplane of the matroid $M|(F\cup A)$. By Lemma 4.6 applied to this matroid with the basis $B$ of $F$, we see that $M$ has a triangle $\{x,y,b\}$ with $x,y\in\operatorname{cl}_{M}(F\cup\{e\})-F$ and $b\in B$, such that $c(x)=\kappa$ and $c(y)=\lambda$.
Since $F\cup A$ is not a decomposer of $M$ by Lemma 7.7(ii), there exists $z\in E(M)-(F\cup A)$ with $c(z)=\lambda$. Let $P=\operatorname{cl}_{M}(\{x,y,z\})$. Since $z\notin F\cup A$ we have $z\in P-\operatorname{cl}_{M}(\{y,b\})$, so $\{y,b,z\}$ is a monochromatic basis of $P$ with colour $\lambda$.
Since $x\in P-H$, the set $H\cup P$ is spanning, so by modularity $r_{M}(H\cap P)=\sqcap_{M}(P,H)=2$, and the flat $H\cap P$ is a line. If $H\cap P\subseteq F$, then $(H\cap P)\cup\{x\}$ is a rank-$3$ subset of $P$ contained in the flat $F\cup A$, and so $P\subseteq F\cup A$, which contradicts $z\in P$. Therefore $H\cap P$ contains at most one element of $F$, and $c((H\cap P)-F)\subseteq c(H-F)=\{\kappa\}$, so (since $M|(H\cap P)$ is connected), the line $H\cap P$ contains at least two elements coloured $\kappa$. Now $x$ is a third element of $P$ coloured $\kappa$ outside this line, so $P$ has a monochromatic basis coloured $\kappa$.
We have just seen that $P$ has a monochromatic basis of each colour. But $(M;c)|P$ is a target by Lemma 7.7 (i), so $P$ has a proper subflat $F_{0}$ with $P-F_{0}$ monochromatic, implying that some colour class of $(M;c)|P$ lies in $F_{0}$ so is nonspanning in $P$. This is a contradiction. ∎
With an additional hypothesis that the lines are not too short, we can upgrade the previous theorem to be stated in terms of lines rather than planes. It is impossible to drop this extra hypothesis, since all rank-$2$ coloured binary projective geometries are targets, but not all coloured binary projective geometries are targets.
The following theorem was also proved for $2$-colourings in [20, Theorem 2]. Our proof is also a little shorter, since it indirectly exploits Theorem 5.9.
Theorem 7.9.
Let $(M;c)$ be a connected modular coloured matroid having a line with at least four points. Then $(M;c)$ is a target if and only if $(M;c)|L$ is a target for each line $L$ of $M$.
Proof.
Let $M$ be a minimum-sized counterexample. By Lemma 7.1, it follows that $M$ is not a target, but $(M;c)|L$ is a target for every line $L$, which implies that $r(M)\geq 3$. By Lemma 7.8, there is a plane $P$ of $M$ such that $(M;c)|P$ is not a target. Since $(M;c)|P$ is a counterexample, we must have $P=E(M)$ by minimality.
By the minimality in the choice of $M$, the coloured matroid $(M;c)|H$ is a target for each hyperplane $H$ of $M$, so $(M;c)$ satisfies the hypotheses of Lemma 7.7. By Lemma 7.7(i), we see that each restriction to a point is a target, so each point is monochromatic. By the minimality in the choice of $M$, it easily follows that $M$ is simple.
Now $M$ is a simple, connected rank-three modular matroid that has a line with at least four elements, so is a projective plane of order $q\geq 3$.
By Lemma 7.7, we have $|c|=2$; let $c(M)=[2]$, and for each $i$ let $X_{i}$ be the set of points of colour $i$. Since $(M;c)|L$ is a target for each line $L$, by Lemma 7.5 we see the following:
7.9.1.
If $i\in[2]$ and $x\neq y\in X_{i}$, then $|X_{i}\cap\operatorname{cl}_{M}(\{x,y\})|\geq q$.
Fix $i\in[2]$. If $X_{i}$ is not spanning, then $E(M)-\operatorname{cl}_{M}(X_{i})$ is nonempty and monochromatic, so $\operatorname{cl}_{M}(X_{i})$ is a decomposer of $(M;c)$, contradicting Lemma 7.7(ii). So $X_{i}$ contains a basis $B_{i}=\{x_{1},x_{2},x_{3}\}$. By 7.9.1, the line $L=\operatorname{cl}_{M}(\{x_{1},x_{2}\})$ satisfies $|L\cap X_{i}|\geq q$, and for each $y\in L\cap X_{i}$, the line $L^{\prime}=\operatorname{cl}_{M}(\{x_{3},y\})$ satisfies $|L^{\prime}\cap X_{i}|\geq q$. The lines in the collection $\mathcal{L}=\{\operatorname{cl}_{M}(\{x_{3},y\}):y\in L\cap X_{i}\}$ pairwise intersect only at $x_{3}$, so
$|X_{i}|\geq 1+\sum_{L^{\prime}\in\mathcal{L}}|(L^{\prime}-x_{3})\cap X_{i}|\geq 1+q(q-1)=q^{2}-q+1.$
But this holds for both $X_{1}$ and $X_{2}$, so $q^{2}+q+1=|M|=|X_{1}|+|X_{2}|\geq 2(q^{2}-q+1)$. This gives $q^{2}+1\leq 3q$, which contradicts $q\geq 3$. ∎
Recall that, for $X\subseteq E(M)$, we write $\omega_{M}(X)$ for the rank of the largest flat of $M$ contained in $X$.
Lemma 7.10.
Let $(M;c)$ be a target for which $M$ is connected and modular, and for each $\lambda\in[c]$, let $X_{\lambda}$ be the colour class of $\lambda$. If $L\subseteq[c]$, then $\sum_{\lambda\in L}\omega_{M}(X_{\lambda})=\omega_{M}(\cup_{\lambda\in L}X_{\lambda})$.
Proof.
Let $F_{1},\dotsc,F_{s}$ be mutually skew monochromatic flats of $M$ with spanning union, such that $c=\bigotimes_{i=1}^{s}(c|F_{i})$, as given by Lemma 7.3. Let $\{\lambda_{i}\}=c(F_{i})$, and let $I=\{i\in\{1,\dotsc,s\}:\lambda_{i}\in L\}$. Then, by Lemma 6.6, we have
$\sum_{\lambda\in L}\omega_{M}(X_{\lambda})=\sum_{\lambda\in L}\sum_{i=1}^{s}\omega_{M}(F_{i}\cap X_{\lambda})=\sum_{i=1}^{s}\sum_{\lambda\in L}\omega_{M}(F_{i}\cap X_{\lambda}).$
Now $F_{i}\cap X_{\lambda}$ is equal to $F_{i}$ if $\lambda=\lambda_{i}$, and is empty otherwise. It follows that $\sum_{\lambda\in L}\omega_{M}(F_{i}\cap\lambda)=r_{M}(F_{i})$ if $i\in I$, and $\sum_{\lambda\in L}\omega_{M}(F_{i}\cap\lambda)=0$ otherwise. Therefore
(1) $\sum_{\lambda\in L}\omega_{M}(X_{\lambda})=\sum_{i\in I}r_{M}(F_{i}).$
Now let $d:[c]\to\{0,1\}$ be the colouring with $d(\lambda)=1$ if and only if $\lambda\in L$, and let $X=(d\circ c)^{-1}(1)$; so $X=\bigcup_{\lambda\in L}X_{\lambda}$ by construction. Since $X$ is a colour class of the colouring $d\circ c=\bigotimes_{i=1}^{s}((d\circ c)|F_{i})$, Lemma 6.6 gives
$\omega_{M}(X)=\sum_{i=1}^{s}\omega_{M}(F_{i}\cap X).$
Now $F_{i}\cap X$ is equal to $F_{i}$ if $i\in I$, and is empty otherwise. So in fact $\omega_{M}(X)=\sum_{i\in I}r_{M}(F_{i})$. Combined with (1), this gives the result. ∎
We now restate and prove Theorem 1.8. In the language of this section, it states that rainbow-triangle-free coloured binary projective geometries where the number of colours is equal to the dimension are all targets, determined by a maximal chain of flats. This implies that they are unique up to matroid isomorphism and permutation of colours.
Theorem 7.11.
Let $q$ be a prime power, and $(M;c)$ be a coloured matroid with no rainbow triangle for which $M\cong\operatorname{PG}(n-1,q)$ and $|c|=n\geq 1$. Then there exist a permutation $\kappa_{1},\dotsc,\kappa_{n}$ of $[c]$ and flats $\varnothing=F_{0}\subset F_{1}\subset\dotsc F_{n}=E(M)$ of $M$, such that $r_{M}(F_{i})=i$ and $c(F_{i}-F_{i-1})=\{\kappa_{i}\}$ for each $i$.
Proof.
It is easy to see, along the lines of the proof of Theorem 7.3, that a coloured matroid $(M;c)$ satisfies the conclusion if and only if $c$ is a lift-join of rank-one monochromatic flats of distinct colours. Thus, it is enough to prove that $M$ has this property.
If $n=2$, then since every triple of points of $M$ is a triangle, the result is straightforward. Suppose inductively that $n\geq 3$, and that the theorem is true for smaller $n$. Since $|c|=n\geq 3$, Theorem 5.9 implies that $(G;c)$ has a nontrivial decomposer $F$.
Let $F^{\prime}$ be a complement of $F$ in $M$. By Lemma 6.3 we have $c=(c|F)\otimes(c|F^{\prime})$. By Theorem 5.2, we have $|c(F)|\leq r_{M}(F)$ and similar for $F^{\prime}$, so
$r_{M}(F)+r_{M}(F^{\prime})=n=|c|=|c(F)|+|c(F^{\prime})|-|c(F)\cap c(F^{\prime})|,$
and so $r_{M}(F)=|c(F)|$ and $r_{M}(F^{\prime})=|c(F^{\prime})|$ while $c(F)\cap c(F^{\prime})=\varnothing$. By induction, both $c|F$ and $c|F^{\prime}$ are lift-joins of monochromatic rank-one flats of distinct colours. The sets of colours are disjoint, and $c=(c|F)\otimes(c|F^{\prime})$, and so the same is true for $M$, as required. ∎
8. Erdős-Hajnal
This section will derive the lower bound in Theorem 1.6 from Theorem 6.8. We first prove that a rainbow-triangle-free coloured matroid can be partioned into two-coloured pieces, in such a way that the partition (viewed as a colouring) is a target. This easily implies Theorem 1.7, since the ‘layers’ of the target $(M,C)$ are bicoloured relative to $c$.
Theorem 8.1.
Let $(M;c)$ be a connected, modular coloured matroid with no rainbow triangle. If $c$ is simple and $|c|\geq 2$, then $M$ has a colouring $C$ for which $C]\subseteq\binom{[c]}{2}$, such that $(M;C)$ is a target, and $c(e)\in C(e)$ for each $e\in E(M)$.
Proof.
The result is vacuously true if $r(M)\leq 1$, since a simple colouring $c$ of $M$ with $|c|\geq 2$ cannot exist. Suppose, therefore, that $r(M)\geq 2$, and that the result holds for coloured matroids of smaller rank.
If $|c|=2$, then we can take $C$ to assign $c]\in\binom{[c]}{2}$ to every element of $M$, and $C$ is a $1$-colouring so is a target. Suppose, therefore, that $|c|\geq 3$.
By Theorem 5.9, $M$ has a nontrivial decomposer $F_{0}$. Let $F_{1}$ be a complement of $F_{0}$ in $M$. Since both $F_{0}$ and $F_{1}$ are proper flats, there inductively exist colourings $C_{0}\subseteq\binom{c(F_{0})}{2}$ and $C_{1}\subseteq\binom{c(F_{1})}{2}$ of $F_{0}$ and $F_{1}$ respectively such that $(M|F_{0};C_{0})$ and $(M|F_{1};C_{1})$ are targets, and $c(e)\in C_{i}(e)$ for all $i\in[2]$ and $e\in F_{i}$.
We claim that the colouring $C_{0}\otimes C_{1}$ works for $M$. First, we have $C_{0}\otimes C_{1}]=[C_{0}]\cup[C_{1}]\subseteq\binom{[c(F_{0})]}{2}\cup\binom{[c(F_{1})]}{2}\subseteq\binom{[c]}{2}$. By Lemma 7.4, we know that $(M;C_{0}\otimes C_{1})$ is a target. It remains to show that $e\in(C_{0}\otimes C_{1})(e)$ for all $e\in E(M)$.
If $e\in F_{0}$, then $(C_{0}\otimes C_{1})(e)=C_{0}(e)\ni c(e)$, as required. If $e\notin F_{0}$, then let $f\in\operatorname{cl}_{M}(F_{0}\cup\{e\})\cap F_{1}$. Since $F_{0}$ is a decomposer of $(M;c)$, we have $c(e)=c(f)$. Since $F_{0}$ is a decomposer of $(M;C_{0}\otimes C_{1})$, we also have $(C_{0}\otimes C_{1})(e)=(C_{0}\otimes C_{1})(f)=C_{1}(f)\ni c(f)=c(e)$, as required. ∎
Since two-coloured pieces in the above theorem form a target, we can easily find a large subspace in one of them. This is analogous to [11, Corollary 1.5].
Theorem 8.2.
Let $(M;c)$ be a connected, modular, rainbow-triangle-free coloured matroid with $c$ simple and $|c|\geq 2$. Then $M$ has a flat $F$ for which $r_{M}(F)\geq r(M)/\binom{|c|}{2}$ and $|c(F)|\leq 2$.
Proof.
Let $C$ be a colouring given by Theorem 8.1, and for each pair $P\in\binom{[c]}{2}$, let $X_{P}$ be the set of elements of $M$ coloured $P$ by $C$. By the choice of $C$, the sets $X_{P}$ partition $E(M)$.
By Lemma 7.10 we have $\sum_{P}\omega_{M}(P)=r(M)$, so there exists $P\in\binom{[c]}{2}$ for which $\omega_{M}(X_{P})\geq r(M)/\binom{|c|}{2}$. Now any flat $F\subseteq C_{P}$ of rank $\omega_{M}(X_{P})$ is two-coloured, so satisfies the theorem. ∎
In fact, the techniques in Theorem 8.2 routinely generalize to the case where we are looking for a large flat using at most some prescribed number of colours. In the following result, Theorem 8.2 is the case where $\ell=2$. More generally, the result is very similar to the upper bound [11, Theorem 1.6].
Theorem 8.3.
Let $\ell,k\in\mathbb{N}$ with $\ell\leq k$ and $k\geq 2$, and let $(M;c)$ be a connected, modular coloured matroid for which $c$ is simple and has no rainbow triangle. If $|c|=k$, then $M$ has a flat $F$ such that $|c(F)|\leq\ell$ and $r_{M}(F)\geq\tfrac{\binom{\ell}{2}}{\binom{k}{2}}r(M)$.
Proof.
We may assume that $\ell\leq k$, as otherwise the result is trivial. Let $C:E(M)\to\binom{[c]}{2}$ be a colouring of $M$ as given by Lemma 8.1, so $(M;C)$ is a target, and $c(e)\in C(e)$ for all $e\in E(M)$.
For each $P\in\binom{[c]}{2}$, let $X_{P}$ be the colour class of $P$ with respect to $C$. Applying Lemma 7.10 with $L=\binom{S}{2}$ and later with $L=\binom{[c]}{2}$, we have
$\sum_{S\in\binom{[c]}{\ell}}\omega_{M}\left(\cup_{P\in\binom{S}{2}}X_{P}\right)$ $=\sum_{S\in\binom{[c]}{\ell}}\sum_{P\in\binom{S}{2}}\omega_{M}(X_{P})$
$=\sum_{P\in\binom{[c]}{2}}\sum_{P\subseteq S\in\binom{[c]}{\ell}}\omega_{M}(X_{P})$
$=\tbinom{k-2}{\ell-2}\sum_{P\in\binom{[c]}{2}}\omega_{M}(X_{P})$
$=\tbinom{k-2}{\ell-2}\omega_{M}\left(\cup_{P\in\binom{[c]}{2}}X_{P}\right)$
$=\tbinom{k-2}{\ell-2}r(M).$
By a majority argument, there is therefore some $S\in\binom{[c]}{\ell}$ for which
$\omega_{M}\left(\cup_{P\in\binom{S}{2}}X_{P}\right)\geq\binom{k}{\ell}^{-1}\binom{k-2}{\ell-2}r(M)=\tfrac{\binom{k}{2}}{\binom{\ell}{2}}r(M).$
By the choice of the colouring $C$ and the $X_{P}$, every element in $\cup_{P\in\binom{S}{2}}X_{P}$ has a colour in the $\ell$-element set $S$, so $\cup_{P\in\binom{S}{2}}X_{P}$ contains a flat $F$ of rank at least $\tfrac{\binom{k}{2}}{\binom{\ell}{2}}r(M)$ with $c(F)\subseteq S$, as required. ∎
We can now prove the lower bound in Theorem 1.6.
Corollary 8.4.
Let $c_{0}$ be a $3$-colouring of $\operatorname{PG}(1,2)$ where the three elements have distinct colours. If $n\geq 2$ and $c$ is a $c_{0}$-free $3$-colouring of $G=\operatorname{PG}(n-1,2)$, then $G$ has a subspace of dimension at least $\tfrac{n}{3}$ containing points of at most two colours.
Proof.
In the connected modular coloured matroid $(M;c)$ corresponding to $G$ and $c$, a rainbow triangle would give a copy of $c_{0}$, so $(M;c)$ is rainbow-triangle-free. By Theorem 8.2, there is a flat $F$ of $M$ with $r_{M}(F)\geq r(M)/\binom{3}{2}=n/3$ and $|c(F)|\leq 2$. This flat gives the required subspace. ∎
9. Lower Bounds
We now prove using standard probabilistic arguments that the bound in Theorem 8.3 (and hence Theorems 8.4 and 8.2) is best-possible up to an additive logarithmic error term. We first need a simple lemma about random red-blue coloured geometries.
Lemma 9.1.
Let $q$ be a prime power, let $n\in\mathbb{N}$, and $\kappa,\lambda$ be colours. If $(M;c)$ is a coloured matroid for which $M\cong\operatorname{PG}(n-1,q)$, and $c$ independently assigns each element of $M$ colour $\kappa$ or $\lambda$ with probability $\tfrac{1}{2}$, then for all $\epsilon>0$, we asymptotically almost surely have $\omega_{M}(X)\leq(1+\epsilon)\log_{q}(n)$ for each colour class $X$.
Proof.
For each $k\in\mathbb{N}$, let $f(k)=\tfrac{q^{k}-1}{q-1}$. Let $i\in\{\kappa,\lambda\}$, and $X=c^{-1}(i)$ be the colour class of $i$. Let $t\in\mathbb{N}$. The number of rank-$t$ flats of $M$ is ${{n}\brack{t}}_{q}$. Each such flat has $f(t)$ elements, so for each $i\in\{\kappa,\lambda\}$, the probability that such a flat is monochromatic with colour $i$ is $2^{-f(t)}$, so by the union bound, the probability that $(M;c)$ has a monochromatic flat with colour $i$ is at most ${{n}\brack{t}}_{q}2^{-f(t)}$. Now
$\log_{2}\left({{n}\brack{t}}_{q}2^{-f(t)}\right)$ $=\sum_{i=1}^{t}\log_{2}\left(\frac{q^{n-i+1}-1}{q^{i}-1}\right)-f(t)$
$\leq\sum_{i=1}^{t}\log_{2}(q^{n})-f(t)$
$\leq tn\log_{2}q-q^{t-1}.$
If $t\geq(1+\epsilon)\log_{q}n$, then $tn\log_{2}q-q^{t-1}\leq(1+\epsilon)n\log_{2}n-q^{-1}n^{1+\epsilon}$. This expression tends to -$\infty$ as $n$ grows, whence we see that almost surely $M$ has no monochromatic flat of rank at least $t$, giving $\omega_{M}(X)\leq t$. ∎
The next result shows that the bound in Theorem 8.3 cannot be improved by more than $2k\log_{q}n$. The true best-possible bound will depend on geometric Ramsey numbers, so is currently out of reach.
Theorem 9.2.
Let $q$ be a prime power, let $\ell,k\in\mathbb{N}$ with $k\geq 2$ and $1\leq\ell\leq k$. For all sufficiently large $n$, there is a rainbow-triangle-free coloured matroid $(M;c)$ with $|c|=k$, such that $M\cong\operatorname{PG}(n-1,q)$, and every flat $F$ of $M$ with $|c(F)|\leq\ell$ satisfies $r_{M}(F)\leq\frac{\binom{\ell}{2}}{\binom{k}{2}}n+2k\log_{q}(n)$.
Proof.
Let $t=\binom{k}{2}$, and let $S_{1},\dotsc,S_{t}$ be an enumeration of $\binom{[c]}{2}$. Let $\epsilon=\tfrac{1}{2(k-\ell)}$, and choose $n$ large enough so that $n\geq q^{\ell}$ and so that Lemma 9.1 applies for $\epsilon,q$, and every matroid of rank at least $\left\lfloor\frac{n}{t}\right\rfloor$.
Let $G_{1},\dotsc,G_{t}$ be a collection of mutually skew flats of $M$ with spanning union, each having rank in $\{\left\lfloor n/t\right\rfloor,\left\lceil n/t\right\rceil\}$. For each $i\in[t]$, let $c_{i}$ be a colouring of $G_{i}$ with the two colours in $S_{i}$ for which there is no monochromatic flat of dimension larger than $\left(1+\epsilon\right)\log_{q}(r_{M}(G_{i}))$, as given by Lemma 9.1. We show that the colouring $c=\bigotimes_{i=1}^{t}c_{i}$ of $M$ satisfies the theorem.
9.2.1.
For each $T\in\binom{[c]}{\ell}$, we have $\omega_{M}(c^{-1}(T))\leq\tfrac{\binom{\ell}{2}}{t}n+\ell k\log_{q}(n)$.
Subproof:.
Let $d:[c]\to\{0,1\}$ be the indicator function of $T$. For each $b\in\{0,1,2\}$, let $I_{b}=\{i\in[t]:|P_{i}\cap T|=b\}$, so $t]=I_{0}\cup I_{1}\cup I_{2}$, and $|I_{b}|=\binom{\ell}{b}\binom{k-\ell}{2-b}$. Let $X=c^{-1}(T)=(d\circ c)^{-1}(1)$. Since $d\circ c=\bigotimes_{i=1}^{t}(d\circ c_{i})$, Lemma 6.6 implies that
$\omega_{M}(X)=\sum_{i=1}^{t}\omega_{M}(G_{i}\cap X)=\sum_{b=0}^{2}\sum_{i\in I_{b}}\omega_{M}(G_{i}\cap X)$
For $i\in I_{0}$, the elements of $G_{i}$ have colour outside $T$, so $G_{i}\cap X=\varnothing$. For $i\in I_{1}$, only one colour assigned to the elements of $G_{i}$ lies in $T$, so $\omega_{M}(G_{i}\cap X)$ is the rank of some monochromatic flat of $c_{i}$, which (by the choice of $c_{i}$) is at most $(1+\epsilon)\log_{q}(r_{M}(G_{i}))\leq(1+\epsilon)\log_{q}(n)$. For $i\in I_{2}$, we have $X\subseteq G_{i}$, so $\omega_{M}(G_{i}\cap X)=r_{M}(G_{i})\leq\tfrac{n}{t}+1$. Combining these estimates gives
$\omega_{M}(X)$ $\leq|I_{2}|(\tfrac{n}{t}+1)+|I_{1}|(1+\epsilon)\log_{q}(n)$
$=\tbinom{\ell}{2}\left(\tfrac{n}{t}+1\right)+\ell(k-\ell)(1+\epsilon)\log_{q}(n)$
$=\tfrac{\binom{\ell}{2}}{t}n+\ell((k-\ell)(1+\tfrac{1}{2(k-\ell)})\log_{q}(n)+\tfrac{1}{2}(\ell-1))$
$\leq\tfrac{\binom{\ell}{2}}{t}n+\ell k\log_{q}(n),$
where the last inequality uses $1\leq\ell\leq\log_{q}(n)$ and the choice of $\epsilon$. ∎
Since $t=\binom{k}{2}$, we have $r_{M}(F)\leq\tfrac{\binom{\ell}{2}}{\binom{k}{2}}n+\ell k\log_{q}(n)$ for every flat $F$, as required. ∎
We state two corollaries showing that the bounds in Theorems 8.2 and 8.4 are nearly sharp, mirroring the lower and upper bounds given in [11, Theorem 1.6]. The second completes the proof of Theorem 1.6.
Corollary 9.3.
Let $q$ be a prime power, and $k\in\mathbb{N}$. For all sufficiently large $n$, there is a rainbow-triangle-free coloured matroid $(M;c)$ with $|c|=k$, such that $M\cong\operatorname{PG}(n-1,q)$, and every flat $F$ of $M$ with $|c(F)|\leq 2$ satisfies $r_{M}(F)\leq n/\binom{c}{2}+2k\log_{q}(n)$.
Corollary 9.4.
For all sufficiently large $n$, there is a rainbow-triangle-free $3$-colouring of $\operatorname{PG}(n-1,2)$, such that every bicoloured subspace has dimension at most $\tfrac{n}{3}+6\log_{2}(n)$.
10. Ternary Lines
We now prove Theorem 1.9, which we rephrase more concretely and in terms of two-colourings rather than induced subgeometries. The key observation is that ternary projective geometries containing no line with exactly two points of each colour are targets by Theorem 7.9.
Theorem 10.1.
Let $c_{0}$ be a $2$-colouring of the four points of $\operatorname{PG}(1,3)$, containing exactly two points of each colour. If $n\geq 2$, and $c$ is a $c_{0}$-free $2$-colouring of $G=\operatorname{PG}(n-1,3)$, then $G$ has a monochromatic subspace of dimension at least $\left\lceil\tfrac{n}{2}\right\rceil$. Moreover, the value of $\left\lceil\tfrac{n}{2}\right\rceil$ is best-possible.
Proof.
If some line $L$ of $M\cong\operatorname{PG}(n-1,3)$ contains exactly two points of each colour, then (since any ordered pair of distinct points of $\operatorname{PG}(1,3)$ is related by a linear isomorphism to any other) there is a copy of $c_{0}$ inside $G$.
It follows that each line of $M$ contains at most one point of some colour, and therefore $(M;c)|L$ is a target for every line $L$. By Theorem 7.9, we see that $(M;c)$ itself is a target. If $X_{1}$ and $X_{2}$ are the colour classes of $c$, we have $\omega_{M}(X_{1})+\omega_{M}(X_{2})=n$, so $\omega(X_{i})\geq\tfrac{n}{2}$ for some $i$, so some $X_{i}$ contains the required monochromatic subspace.
To see that $\left\lceil\tfrac{n}{2}\right\rceil$ is best possible, consider a $2$-colouring $c$ of the matroid $M=\operatorname{PG}(n-1,3)$ in which one colour class is a rank-$\left\lceil\tfrac{n}{2}\right\rceil$-dimensional flat $F$. Every monochromatic flat $F^{\prime}$ of $M$ with rank greater than $\tfrac{n}{2}$ must intersect $F$ due to a dimension argument, so must be contained in $F$ by the construction of $c$; but then $\left\lceil\tfrac{n}{2}\right\rceil<r_{M}(F^{\prime})\leq r_{M}(F)=\left\lceil\tfrac{n}{2}\right\rceil$, a contradiction. ∎
11. Claws
The following was proved in [6], Lemma 4.5. There, its statement is phrased using the ‘critical number’ $\chi$, which for a set $X$ of elements in a matroid $M\cong\operatorname{PG}(n-1,2)$ is defined by $\chi(X)=n-\omega_{M}(E(M)-X)$.
Lemma 11.1.
If $M\cong\operatorname{PG}(n-1,2)$ and $X\subseteq E(M)$ is a set such that $|X\cap P|$ is even for every plane $P$ of $M$, then $\omega_{M}(E(M)-X)\geq\left\lceil\tfrac{n}{2}\right\rceil-1$.
The next lemma follows from Lemma 3.3.
Lemma 11.2.
[…]
Proof.
[…]
For the next two results, let $\widehat{c}$ be a $2$-colouring of the binary projective plane $\operatorname{PG}(2,2)$ in which the colour class of $1$ is a three-element independent set. Viewed as a binary matroid rather than a colouring, this is the claw. The main theorem of [21], which we phrase here in terms of colourings, precisely describes the structure of claw-free binary matroids.
Theorem 11.3.
Let $c$ be a $\widehat{c}-$free 22-colouring of $M\cong\operatorname{PG}(n-1,2)$ with colour classes $X_{1}$ and $X_{2}$. Then there are mutually skew flats $F_{1},\dotsc,F_{t}$ of $M$ such that $c=\bigotimes_{i}(c|F_{i})$, and for each $i$, either
  1. (i)
    $X_{1}\cap F_{i}$ is the union of two disjoint flats of $F_{i}$ whose union is spanning in $F_{i}$,
  2. (ii)
    $\omega_{M}(X_{2}\cap F_{i})\leq 1$, or
  3. (iii)
    for each plane $P$ of $F_{i}$, the set $X_{1}\cap P$ has even size.
We can now prove Theorem 1.10, which we rephrase in terms of two-colourings.
Theorem 11.4.
There exists $C>0$ such that, if $c$ is a $\widehat{c}-$free two-colouring of $M\cong\operatorname{PG}(n-1,2)$, then $M$ has a monochromatic flat of rank at least $Cn^{1/7}$.
Proof.
[…]
For each $i\in I_{2}$, applying Lemma 11.2 to $F_{i}\cap X_{1}$ in the matroid $M|F_{i}$ gives $\omega_{M}(F_{i}\cap X_{1})\geq C^{\prime}(r_{M}(F_{i}))^{1/7}$. It follows from Lemma 6.6 that
$\omega_{M}(X_{1})\geq\sum_{i\in I_{1}\cup I_{2}}C^{\prime}(r_{M}(F_{i}))^{1/7}\geq C^{\prime}\left(\sum_{i\in I_{1}\cup I_{2}}r_{M}(F_{i})\right)^{1/7}.$
For each $i\in I_{3}$, Lemma 11.1 gives $\omega_{M}(F_{i}\cap X_{2})\geq\tfrac{1}{2}r_{M}(X_{i}),$ so
$\omega_{M}(X_{2})\geq\sum_{i\in I_{3}}\tfrac{1}{2}r_{M}(F_{i})=\tfrac{1}{2}\sum_{i\in I_{3}}r_{M}(F_{i}).$
Since the $F_{i}$ are skew with spanning union, we have $\sum_{i}r_{M}(F_{i})=n$, so either $\sum_{i\in I_{1}\cup I_{2}}r_{M}(F_{i})\geq\tfrac{n}{2}$ or $\sum_{i\in I_{3}}r_{M}(F_{i})\geq\tfrac{n}{2}$. The first implies that $\omega_{M}(X_{1})\geq(C^{\prime}2^{-1/7})n^{1/7}=Cn^{1/7}$ and the second gives $\omega_{M}(X_{2})\geq\tfrac{n}{2}\geq Cn^{1/7}$, as required. ∎
12. Infinite Fields
Matroids are traditionally defined to have finite ground sets, but this is certainly not necessary. Infinite matroids of finite rank can easily be defined from the usual matroid rank axioms by simply dropping the requirement that the ground set is finite while maintaining the stipulation that rank is integer-valued. With the notable exception of matroid duality, most matroid theory works exactly as expected for these objects – see [5] for a discussion. It is natural to ask whether Theorem 6.8 holds in this expanded setting, since it would then apply to colourings of finite-dimensional projective space over fields such as $\mathbb{Q},\mathbb{R}$ and $\mathbb{C}$. Indeed, every component of the proof of Theorem 6.8 easily adapts to finite-rank infinite matroids, except Lemma 5.7, whose proof involves counting points in lines.
Unfortunately, this issue is fundamental, because the generalization of Theorem 6.8 to projective geometries over infinite fields turns out to be false. We show via an algebraic construction that for many infinite fields $\mathbb{F}$ (including $\mathbb{Q},\mathbb{R}$ and $\mathbb{C}$), there is a rainbow-triangle-free $n$-colouring of $\operatorname{PG}(n-1,\mathbb{F})$ that cannot be decomposed via lift-joins. The construction is a higher-dimensional version of one discovered for projective planes by Carter and Vogt [7] and, independently, Hales and Straus [14].
To explain the construction, we need some algebra. A non-Archimedean valuation on a field $\mathbb{F}$ is a function $v:\mathbb{F}\to\mathbb{R}\cup\{\infty\}$ so that $v^{-1}(\infty)=\{0\}$, and $v(ab)=v(a)+v(b)$ and $v(a+b)\geq\min(v(a),v(b))$ for all $a,b\in\mathbb{F}$. Note that $v(a^{-1})=-v(a)$ for all $a$ and therefore that v(1)=v(1)=0v(1)=v(-1)=0; if there is some $a\neq 0$ for which $v(a)\neq 0$, then $v$ is nontrivial.
It is routine to show that finite fields do not admit nontrivial non-Archimedean valuations, and it is known that the same is true for any algebraic extension of a finite field. Many other fields do have such valuations, however. For example, if $p$ is a prime, then the $p$-adic valuation on $\mathbb{Q}$, defined by $v(p^{t}\tfrac{r}{s})=t$ for all $r,s$ relatively prime to $p$, is non-Archimedean. Nontrivial valuations on $\mathbb{R}$ and $\mathbb{C}$ are harder to construct, but their existence follows from the axiom of choice.
[…]
It is easy to check using the definition of a non-Archimedean valuation that $c_{v}(\lambda u)=c_{v}(u)$ for all $\lambda\in\mathbb{F}^{\times}$, so $c_{v}$ can also be viewed as an $n$-colouring defined on the projective space $\operatorname{PG}(n-1,\mathbb{F})$ whose points are the one-dimensional subspaces of $\mathbb{F}^{n}$.
(For $p$-adic valuations on the field of rationals, we can view $c_{v}$ very concretely: to colour a point $x=(r_{1},\dotsc,r_{n})$, simply scale the $r_{i}$ to be integers with no common factor, and colour $x$ with the first index whose entry is not a multiple of $p$.)
[…]
We first show that the colouring $c_{v}$ is rainbow-triangle-free – in fact it is just as easy to show the equivalent statement that it has no rainbow circuit. (Showing this for all circuits has the tiny advantage of not exploiting the equivalence in Lemma 5.1, which does hold for finite-rank infinite matroids, but we technically did not prove in this setting.)
Theorem 12.1.
Let $n\geq 1$, and let $v$ be a non-Archimedean valuation on a field $\mathbb{F}$. Then every $c_{v}$-rainbow set in $\operatorname{PG}(n-1,\mathbb{F})$ is independent.
Proof.
Let $I=\{x^{1},\dotsc,x^{t}\}$ be a $c_{v}$-rainbow set, where $c_{v}(x^{1})<c_{v}(x^{2})<\dotsc<c_{v}(x^{t})$. Let $x^{i}=[(x^{i}_{1},\dotsc,x^{i}_{n})]$ and $c_{v}(x^{i})=k_{i}$ for each $i$.
Write $|x|$ for $v(x)$. Since $c_{v}(x^{i}_{1},\dotsc,x^{i}_{n})=k_{i}$ for each $i$, we have $|x^{i}_{k_{i}}|\leq|x^{i}_{j}|$ for all $j$, and the inequality is strict if $j<k_{i}$.
To show that $I$ is independent, it suffices to show that the $t\times t$ matrix $A$ defined by $A_{ij}=x^{i}_{k_{j}}$ is nonsingular. Suppose, therefore, that the determinant $|A|$ of $A$ is zero. Write $s(\pi)=\pm 1$ for the sign of a permutation $\pi\in S_{n}$. For each $\pi\in S_{n}$, using $|s(\pi)|=0$, we have
$\left|s(\pi)\prod_{i=1}^{t}x^{i}_{k_{\pi(i)}}\right|=\sum_{i=1}^{t}|x^{i}_{k_{\pi(i)}}|\geq\sum_{i=1}^{t}|x^{i}_{k_{i}}|.$
Moreover, if $\pi$ is not the identity permutation, there is some $s$ such that $\pi(s)<s$, which implies that $k_{\pi(s)}<k_{s}$, so $|x^{s}_{k_{\pi(s)}}|>|x^{s}_{k_{s}}|$, and the inequality above is strict. Therefore
$\sum_{i=1}^{t}|x^{i}_{k_{i}}|$ $<\min_{\pi\neq 1}\left|s(\pi)\prod_{i=1}^{t}x^{i}_{k_{\pi(i)}}\right|\leq\left|\sum_{\pi\neq 1}s(\pi)\prod_{i=1}^{t}x^{i}_{k_{\pi(i)}}\right|=\left||A|-\prod_{i=1}^{t}x^{i}_{k_{i}}\right|.$
But $|A|=0$ and $\left|-\prod_{i=1}^{t}x^{i}_{k_{i}}\right|=\sum_{i=1}^{t}|x^{i}_{k_{i}}|$, giving a contradiction. ∎
If $F$ is a nontrivial decomposer of a coloured projective geometry $(G;c)$ with $|c|\geq 2$, there clearly exist $x\in F$ and $y\neq F$ with $c(x)\neq c(y)$, and $x$ is the only point of the line $\operatorname{cl}_{M}(\{x,y\})$ with colour $c(x)$.
The next theorem, on the other hand, shows that colourings $c_{v}$ arising from nontrivial non-Archimedean valuations contain no lines that intersect a colour class in exactly one point, and hence they cannot be decomposed via lift-joins. This in turn implies that the statement of Theorem 6.8 must fail for projective geometries over general infinite fields, even when it is adjusted to allow for infinite matroids.
Theorem 12.2.
Let $n\geq 1$, and $v$ be a nontrivial non-Archimedean valuation on a field $\mathbb{F}$. Then no line $L$ of $\operatorname{PG}(n-1,\mathbb{F})$ intersects a colour class of $c_{v}$ in exactly one point.
Proof.
Suppose that $x=[(x_{1},\dotsc,x_{n})]$ is the unique point of a line $L$ having colour $k\in[n]$. Since $L$ contains no linearly independent triple, it contains no rainbow triple by Theorem 12.1, so there is some colour $\ell\neq k$ so that $c(L-\{x\})=\{k\}$.
Since $v$ is nontrivial, there exists $d\in\mathbb{F}$ with $v(d)<0$. Define hyperplanes $W_{1}=\{a\in\mathbb{F}^{n}:a_{k}=0\}$ and $W_{2}=\{a\in\mathbb{F}^{n}:a_{\ell}=da_{k}\}$ of $\mathbb{F}^{n}$, and let $H_{1},H_{2}$ be the corresponding hyperplanes of $\operatorname{PG}(n-1,\mathbb{F})$.
For each $a\in W_{1}$, we have $a_{k}=0$ and so $v(a_{k})=\infty$ and hence $c_{v}(a)\neq k$ by the definition of $c_{v}$. Similarly, for each $a\in W_{2}$, we have $v(a_{\ell})=v(d)+v(a_{k})<v(a_{k})$, so $c_{v}(a)\neq k$. But since $L$ is a line, there are points $f_{1}\in L\cap H_{1}$ and $f_{2}\in L\cap H_{2}$. As just observed, we have $c_{v}(f_{1})\neq k$ and $c_{v}(f_{2})\neq k$; since $f_{1},f_{2}\in L$ we must have $f_{1}=f_{2}=x$.
Therefore $x\in H_{1}\cap H_{2}$, which by the definitions of $H_{1}$ and $H_{2}$ implies that $x_{k}=x_{\ell}=0$. This contradicts the fact that $c_{v}(x)=k$. ∎
On the other hand, algebraic extensions of finite fields do not have nontrivial non-Archimedean valuations, and Theorem 1 of [14] will imply Lemma 5.7 for projective planes over all such fields. Using this fact, one could routinely extend Theorem 6.8 to apply to all finite-dimensional coloured projective geometries over algebraic extensions of finite fields.
Acknowledgements
This project was initiated at the Workshop on (Mostly) Matroids held at the Institute for Basic Science in Daejeon, South Korea in 2024. We thank Rutger Campbell and the other organizers of the workshop for providing an extremely stimulating research environment. We also thank Denys Bulavka for helpful conversations, and Sean McGuinness for alerting us to the related work in [2]. Tony Huynh is supported by the Institute for Basic Science (IBS-R029-C1).